### v1

# Improving Factuality for Dialogue Response Generation via Graph-Based Knowledge Augmentation

Abstract
Large Language Models (LLMs) succeed in many natural language processing tasks. However, their tendency to hallucinate — generate plausible but inconsistent or factually incorrect text — can cause problems in certain tasks, including response generation in dialogue. To mitigate this issue, knowledge-augmented methods have shown promise in reducing hallucinations. Here, we introduce a novel framework designed to enhance the factuality of dialogue response generation, as well as an approach to evaluate dialogue factual accuracy. Our framework combines a knowledge triple retriever, a dialogue rewrite, and knowledge-enhanced response generation to produce more accurate and grounded dialogue responses. To further evaluate generated responses, we propose a revised fact score that addresses the limitations of existing fact-score methods in dialogue settings, providing a more reliable assessment of factual consistency. We evaluate our methods using different baselines on the OpendialKG and HybriDialogue datasets. Our methods significantly improve factuality compared to other graph knowledge-augmentation baselines, including the state-of-the-art G-retriever. The code will be released on GitHub  [^1] https://github.com/XiangyanChen/Triton.
Improving Factuality for Dialogue Response Generation via Graph-Based Knowledge Augmentation
Xiangyan Chen1  Yujian Gan1  Matthew Purver1,2 1Queen Mary University of London, UK 2Institut Jožef Stefan, Slovenia {xiangyan.chen, y.gan, m.purver}@qmul.ac.uk
1 Introduction
Figure 1: Examples comparing an LLM’s dialogue response generation when utilising retrieved triples and a rewritten dialogue versus generating responses without external knowledge or dialogue rewriting. The knowledge triple retriever module assists in selecting the triples most relevant to the query.
Large Language Models (LLMs) have been shown to perform powerfully on Natural Language Processing (NLP) tasks. Despite their general superiority, LLMs will generate some plausible but fact-inconsistent text, namely hallucination. Flawed pre-training data, model bias and randomness in inference are the factors contributing to hallucinations Zhang et al. (2023); Huang et al. (2023). In dialogue response generation, generating incorrect responses will mislead people and have a negative impact on society.
A number of methods have been proposed to enhance the factuality of language models. Among them, knowledge-augmented inference methods aim to do this by adding knowledge triples into the prompt. This has been effective in tasks such as question-answering Baek et al. (2023); Sen et al. (2023); Wu et al. (2023).
However, these works Baek et al. (2023); Wu et al. (2023) tend to use simple sentence embedders to encode knowledge and determine the match between a query and knowledge based on similarity, rather than employing a specialised model designed to assess the relevance between the query and knowledge. He et al. (2024) proposed a graph-based method called G-retriever, which employs the Prize-Collecting Steiner Tree to help select knowledge triples. However, applying these methods to dialogue response generation is challenging due to their lack of consideration for dialogue context. Dialogues frequently involve intricate coreference structures, which complicate entity linking and hinder the LLMs’ comprehension, ultimately compromising the factual consistency and quality of the generated responses.
To address above limitations, we propose a novel framework to improve dialogue factuality that consists of three key components: a knowledge triple retriever, a dialogue rewriting and knowledge-enhanced dialogue response generation. The knowledge triple retriever is fine-tuned on our well-collected samples, which helps select the valuable triples for the dialogue. The dialogue rewrite is based on Chain of Thought (CoT), which resolves the dialogue coreference. The knowledge-enhanced dialogue generation is proposed to improve the dialogue factuality, based on two kinds of approaches: prompt-based and graph-based. Figure 1 illustrates how the knowledge-augmented generation works. The query asks about what movies star “Diane Lane”, and the initial response generated by the LLM contains factual inaccuracies. After applying a dialogue rewrite module that resolves coreference, a triple retriever helps select the most relevant N triples, including (“Nights in Rodanthe”, “cast member”, “Diane Lane”). With these corrected triples, the model generates an accurate response.
To evaluate factuality for responses, we primarily rely on the fact score Min et al. (2023). However, the original design of the fact score does not consider dialogue context and situations where the knowledge source is unavailable, making it difficult to evaluate dialogue responses. To address this limitation, we adapt the fact score to the dialogue settings and assess model-human agreement. Annotation results show that our adapted fact score achieves substantial model-human agreement. Additionally, we propose Not Enough Information Proportion (NEIP) to evaluate dialogue factuality comprehensively. It is a metric that quantifies the proportion of atomic facts in a response that cannot be verified, such as opinions or hallucinated content that can not find any evidence to verify. We utilize the F1 score as a supplementary metric to assess factual accuracy and evaluate response quality through BLEU Reiter (2018), ROUGE-L Lin (2004), and Perplexity (PPL) Jelinek et al. (1977). Additionally, we perform human evaluations to further assess response quality.
We compare our proposed method not only with standalone LLMs but also with knowledge-enhanced approaches such as KAPING, the BM25 algorithm, and the State-of-the-Art (SOTA) G-retriever He et al. (2024). Experimental results show that our framework consistently outperforms both existing knowledge-enhanced methods and the current best-performing G-retriever.
Our contributions to this work can be listed as:
  1. 1.
    We propose a novel framework designed to enhance the factual accuracy of dialogue response generation. This framework incorporates three key components: a knowledge triple retriever that selects valuable triples, a dialogue rewrite that resolves coreference, and a knowledge-enhanced mechanism that effectively boosts overall factuality.
  2. 2.
    We adapt the fact score, originally designed for biography generation, to evaluate the factuality of dialogue systems, ensuring its validity within our task. The evaluation is fine-grained, enabling a more reliable assessment of factual consistency in dialogue systems.
  3. 3.
    We validate our methods against various baselines using two public dialogue datasets. Experimental results demonstrate that our approach significantly outperforms existing methods in factuality while keeping good quality at the generated response.
2 Related Work
2.1 LLM Hallucinations and Mitigation Methods
LLMs are trending in NLP, whose architecture can mainly be classified as encoder-only Devlin (2018), decoder-only Brown (2020); Dubey et al. (2024), and encoder-decoder Chung et al. (2024). However, hallucination exists in these LLMs widely, which is the phenomenon that language models generate coherent but fact-inconsistent text. Many reasons contribute to hallucination. Huang et al. (2023) declared one potential reason is that the data in the pre-training of LLMs is incomplete, incorrect or outdated. Other reasons for hallucinations are the bias of LLMs and randomness in inference. Zhang et al. (2023) categorise hallucinations as input-conflicting, context-conflicting and fact-conflicting, with recent focus on the latter.
Several knowledge-based methods have been proposed to mitigate fact-conflicting hallucinations Agrawal et al. (2023), which generally fall into three categories: knowledge-aware inference, knowledge-aware training, and knowledge-aware validation. For instance, Baek et al. (2023) introduced KAPING, a prompt-based framework for question answering that retrieves knowledge triples based on embedding similarity. Building upon retrieval techniques, Sen et al. (2023) combined Knowledge Graph (KG) retrieval with language model reasoning to enhance performance on complex questions. Similarly, CRAG Yan et al. (2024) employs an evaluator to assess generation quality and, if necessary, refines outputs via web search, showing significant improvements across various generation tasks. Another notable approach is Self-RAG Asai et al. (2023), which retrieves relevant knowledge, generates an initial answer, and then refines it through self-critique, achieving strong results in multiple question answering benchmarks. In a related development, He et al. (2024) proposed G-Retriever, which integrates a graph-based encoder with the Prize-Collecting Steiner Tree algorithm to encode retrieved graph knowledge effectively, standing out among graph-enhanced knowledge retrieval methods.
2.2 Evaluation of Factuality
Dialogue response generation differs from question answering in that it must not only take the dialogue context into account when generating responses, but also must be evaluated based on the factuality and quality of the responses, rather than just the accuracy of answers. Previous dialogue factuality evaluation mainly focuses on humans Ni et al. (2023); Li et al. (2022); Yu et al. (2022), and it is inefficient.
Automatic evaluating factuality is challenging, however; current methods are either reference-based or reference-free.
Several question-answering datasets provide factual references in the form of entities; reference-based evaluation can then be based directly on entity matching. Other NLP datasets offer text references, so an alternative reference-based method involves matching the extracted entities between the generated text and these references Nan et al. (2021). Instead, dialogue datasets provide reference responses. To evaluate the factual accuracy of the generated text, we can therefore compare the entities extracted from the generated response with those in the reference response, measured as F1 score. However, relying solely on this can be inadequate, as the reference may not always provide explicit answers. It is crucial to consider additional metrics to ensure a comprehensive evaluation.
Reference-free methods can be categorised into uncertainty estimation Farquhar et al. (2024) and external knowledge-based approaches, with the former often limited by its reliance on the generation model’s own confidence. The fact score Min et al. (2023), which falls into the latter category, is a metric designed to measure fact consistency in long-form text. The process begins with an LLM breaking down the text into fine-grained sentences called atomic facts. These atomic facts are then verified using both the LLM and external knowledge sources (Min et al. (2023) use Wikipedia titles). The fact score is calculated based on the precision.
However, the fact score is not directly suitable for our work for two key reasons: First, it evaluates only the generated text without taking the dialogue history into account, which is crucial when assessing the factuality of dialogue responses. Second, it does not account for the possibility that generated responses may lack external knowledge sources for verification.
Figure 2: The workflow of dialogue response generation and factuality evaluation. The left part of the figure is our framework. It starts with rewriting the dialogue, and then the knowledge triples will be retrieved from KG. The triple retriever helps select valuable triples. The colour-coded arrows marked as Serial Number 6 represent different knowledge-enhanced methods: the orange arrow indicates the graph-based approach, while the blue arrow corresponds to the prompt-based approach. The right part is factuality evaluation, in which the response will be broken down into individual and verifiable facts, checked using both the LLM and external knowledge sources.
3 Methodology
The diagram in Figure 2 illustrates the workflow of dialogue generation (left side) and factuality evaluation (right side).
Our proposed framework consists of three key components: (1) A well-designed knowledge triple retriever and see Section 3.2 for details, (2) A CoT-based dialogue rewriter, described in Section 3.3, and (3) the knowledge-enhanced dialogue response generation (Section 3.2).
For factuality evaluation (Section 3.5), we adopt the fact score as a primary metric. However, as the original fact score does not fully capture certain aspects of dialogue evaluation, we introduce modifications to better align with our objectives.
3.1 Task Formulation
Our task is to generate a response $s$ given a dialogue context $c$, query $x$ and set of triples $f=\{(h_{1},r_{1},t_{1}),(h_{2},r_{2},t_{2}),...,(h_{n},r_{n},t_{n})\}$ from an encyclopedia-based KG, where $h$ and $t$ represent the head and tail entities, and $r$ denotes the relation between them.
3.2 Knowledge Retriever
The KG is a collection of triples, and these triples are the form of head entity, relation, and tail entity. In encyclopedia-based KGs, like Wikidata Vrandečić and Krötzsch (2014), a triple usually indicates a fact. We aim to retrieve triples from the KG for the query in dialogue and incorporate triples into prompts as a supplement.
We first extract entities from the query using the entity linking approach, and then, we match all entities with the triples from the KG when the entity equals the head or tail entity. Since our entities are sourced from Wikidata, and ReFinED Ayoola et al. (2022) is an entity linking method based on corresponding Wikidata entities, we adopt it for our retrieval process.
Knowledge retriever gathers all triples associated with the entities extracted from the query, but not all of these are relevant to the query. Irrelevant triples introduce textual noise, which has a negative impact on text generation. Thus, it is crucial to select the most relevant triples for the query. Traditional methods rank triples based on word frequency using BM25 Robertson et al. (2009), which primarily captures word overlap. Alternatively, sentence embedding models like MPNet Song et al. (2020), as employed by KAPING Baek et al. (2023), encode semantic information but in a relatively straightforward manner.
To retrieve delicate knowledge, we collect query-triple pairs to fine-tune LLM as a knowledge matcher.
Data Collection
Manual data annotation is both time-consuming and labour-intensive. Given GPT’s strong performance in NLP tasks, we adopt it as an annotation tool. Our approach extracts query-triple pairs from the training dataset and prompts GPT-4o to determine their relevance. The model classifies each pair as either relevant or irrelevant. Based on our manual assessment, GPT-4o demonstrates high accuracy in determining query-triple relevance, described in Appendix F.
Fine-Tuning
After collecting these data, we fine-tune an LLM as the binary classifier. In our work, we select the Llama3 8B model Dubey et al. (2024). We combined LoRA Hu et al. (2021) with fine-tuning, which is an efficient fine-tuning method that adopts low-rank modules into transformer layers. Cross-entropy loss is used to optimize the model during fine-tuning, and it is formulated as follows:
$\mathcal{L}=-\sum_{i=1}^{N}y_{i}\log(p_{i})$ (1)
where $y_{i}$ denotes the true label, $p_{i}$ is defined as the predicted probability and $N$ is the number of label classifications.
With the fine-tuned LLM $\mathcal{M}_{t}$, the query $x$ and the triple $f_{i}$, the probability of relevance can be defined as:
$P(y=\text{relevant}|x,f_{i})=\mathcal{M}_{\text{t}}(x,f_{i})$ (2)
Then, the selected triples are formulated as:
$f_{\text{sorted}}=\text{sort}_{\text{desc}}(P,f,x)$ (3)
where $\text{sort}_{\text{desc}}$ denotes the sort function descending by $P$, and $P$ is the relevance probability distribution.
3.3 Dialogue Rewrite
Unlike in question answering tasks, dialogue response generation involves coreferences that pose challenges for both entity linking and response generation. An intuitive approach to improve entity linking is to expand the scope for extracting entities from the dialogue context. However, this approach may also introduce irrelevant entities, negatively impacting performance.
To resolve coreference and improve the performance of entity linking and dialogue response generation, we propose a CoT-based coreference resolution.
Data Collection
We prompt GPT-4o with dialogue to resolve the conversation in a zero-shot CoT setting, as described by Kojima et al. (2022). Step-by-step reasoning in LLMs enhances the generation of accurate results and aids in fine-tuning smaller language models.
Fine-Tuning
We fine-tune the Llama3 8B model using LoRA on collected coreference resolution samples. The reason for selecting LoRA for this task is that it will not destroy the LLMs reasoning capability obtained from pre-training, making it perform better on this task. Cross-entropy loss is also selected for this task.
After fine-tuning, we prompt LLM to resolve coreference, and the dialogue is rewritten. The rewrite and corresponding sorted triples are formulated as follows:
$c^{\prime},x^{\prime}=\mathcal{M}_{\text{f}}(p_{\text{rewrite}},c,x)$ (4)
$f_{\text{sorted}}^{{}^{\prime}}=\text{sort}_{\text{desc}}(P,f,x^{\prime})$ (5)
where $p_{\text{rewrite}}$, described in Table 6 of Appendix A, is the prompt for rewriting dialogue. $\mathcal{M}_{\text{f}}$ is the LLM for dialogue rewriting.
3.4 Knowledge-Based Dialogue Response Generation
To mitigate hallucinations of LLM and have a factual response, we offer two patterns of generating dialogue responses given selected knowledge triples: prompt-based and graph-based methods.
Prompt Based Generation
Given dialogue context $c$, query $x$, rewritten dialogue $c^{\prime}$ and $x^{\prime}$, as well as selected triples $f_{sorted}$ and $f_{sorted}^{{}^{\prime}}$, the generation of knowledge-augmented dialogue responses and their counterparts with the rewritten dialogue are formulated as follows:
$s=\mathcal{M}_{\text{g}}(p_{\text{gen}},c,x,f_{\text{sorted}})$ (6)
$s^{{}^{\prime}}=\mathcal{M}_{\text{g}}(p_{\text{gen}},c^{\prime},x^{\prime},f_{\text{sorted}}^{{}^{\prime}})$ (7)
where $p_{\text{gen}}$ denotes the prompt for generating the response. $\mathcal{M}_{\text{g}}$ is the LLM for response generation. Table 6 in Appendix A describes the details of the prompt template.
Graph-Based Generation
Prompt-based generation can effectively improve factuality with training, while it is unable to capture the connectivity information within a Knowledge Graph. To address this limitation, we encode the knowledge graph with GNN (Graph Netural Network). We first convert the sorted triples $f^{{}^{\prime}}_{sorted}$ into a graph $\mathcal{G}$. Then the graph and dialogue encoding are defined as follows:
$v=E_{graph}(\mathcal{G})$ (8)
$z=\mathcal{M}(c^{\prime},x^{\prime})$ (9)
where $E_{graph}$ is GNN. $\mathcal{M}$ represents an arbitrary LLM. $v$ and $z$ denote the graph and text embedding, respectively.
We generate the dialogue response by decoding the concatenation of the graph and text embedding, formulated as follows:
$s_{g}=Decoder([v;z])$ (10)
where ; denotes concatenation and $s_{g}$ is the response generated by graph-based method. To preserve the capability of the LLM, we combine LoRA with cross-entropy to train our graph model.
3.5 Evaluation Metrics on Fact
Fact score Min et al. (2023) measures the fact consistency for long-form text. As we mentioned in Section 2.2, it does not consider the dialogue history and the possibility that the generated response may lack the knowledge source to support it.
The atomic facts and fact verification prompts of fact score are modified in this paper. In the revised atomic fact-splitting prompt, only complete sentences are permitted to be split, and the model is prohibited from adding any additional information to the atomic facts. We incorporate dialogue history and introduce the “Not Enough Information” label for the revised fact verification prompt. If an atomic fact is not a factual claim or lacks direct support from any known sources, the model will output ‘Not Enough Information”. The detailed prompts for atomic fact-splitting and fact verification are provided in Table 8 of Appendix A.
Since the original fact score framework does not specify how to get knowledge sources for dialogue, we extract entities from atomic facts and query as Wikipedia titles to search related passages as knowledge sources.
4 Experiment
4.1 Dataset
Two public dialogue datasets, OpendialKG Moon et al. (2019) and HybriDialogue Nakamura et al. (2022), are used to evaluate our methods. The two datasets both provide response references but not factual ones. See Appendix B for the details and examples of datasets.
4.2 Conventional Evaluation Metrics
We aim to evaluate factuality and quality for dialogue response generation. BLEU Reiter (2018), ROUGE-L Lin (2004), Perplexity Jelinek et al. (1977) and F1 score Nan et al. (2021) are employed to evaluate the quality and fact consistency of the generated responses. For more details, please refer to Appendix E.
4.3 Modified Evaluation Metrics
We adopt both the adapted fact score and the NEIP (Not Enough Info Proportion) to comprehensively evaluate factuality. We also manually assess fact score with humans, as shown in Table 1. See the Appendix F for more annotation details.
Datasets Agreement Cohen’s Kappa
HybriDialogue 0.78 0.65
OpendialKG 0.78 0.67
Table 1: The agreement and Cohen’s Kappa between the ground truth and LLM outputs regarding the fact score.
[…]
ChatGLM-6B 1.42 14.24 12.98 70.91 50.13 15.14
   +BM25 1.38 14.02 13.15 70.99 49.07 14.92
   +KAPING 1.31 13.93 13.15 72.37 48.74 15.34
   +Triton((Ours) 1.37 14.07 13.31 74.64 48.73 15.40
   +Triton-C(Ours) 1.34 13.47 14.41 76.46 37.93 15.33
[…]
   +BM25 2.65 12.90 18.92 69.32 56.09 16.25
   +KAPING 2.56 12.72 18.18 68.34 55.79 16.03
   +Triton(Ours) 2.60 12.66 18.89 71.56 54.28 16.09
   +Triton-C(Ours) 2.19 11.27 20.22 74.25 28.53 15.00
[…]
   +BM25 2.68 12.71 23.18 72.06 31.98 13.69
   +KAPING 2.91 12.82 23.22 73.00 32.07 13.85
   +Triton(Ours) 3.16 13.06 23.07 76.20 31.42 14.50
   +Triton-C(Ours) 2.33 10.69 20.03 77.77 20.35 13.78
[…]
Triton-X(Ours) 3.99 25.86 14.43 87.73 59.20 13.75
Table 2: The experimental result with various methods on the OpendialKG dataset. The column heads with * indicate our primary metrics for evaluating factual accuracy. Ours means generating responses with our proposed methods. The bold number is the best result within each model. PPL indicates Perplexity and a lower PPL means more fluency. NEIP denotes the “not enough information” proportion, representing the percentage of atomic facts that either lack direct knowledge support or fail to qualify as factual claims.
4.4 Baseline Methods
Several widely recognised LLMs, such as ChatGLM Zeng et al. (2022) and Flan-T5 Chung et al. (2022), have been selected for this work. Furthermore, we compare our proposed Triton, Triton-C and Triton-X with BM25 Robertson et al. (2009), KAPING Baek et al. (2023), G-Retriever He et al. (2024), and a basic dialogue generation model. Additional details of LLMs can be found in the Appendix G.
[…]
ChatGLM-6B 4.62 24.03 23.66 46.58 58.19 39.57
   +BM25 4.68 24.19 23.77 53.32 55.18 40.26
   +KAPING 4.56 24.07 23.58 55.84 54.51 40.99
   +Triton(Ours) 4.71 24.43 23.95 58.15 53.07 40.96
   +Triton-C(Ours) 4.53 24.23 26.03 59.10 50.99 40.92
[…]
   +BM25 8.71 25.81 32.79 67.42 47.43 38.50
   +KAPING 8.92 26.43 31.96 68.92 46.56 40.41
   +Triton(Ours) 8.90 26.14 30.81 70.79 43.52 39.70
   +Triton-C(Ours) 8.20 24.63 34.01 74.41 33.68 40.63
[…]
   +BM25 11.08 29.75 38.89 60.40 48.79 44.65
   +KAPING 10.95 29.44 37.88 63.84 48.94 44.13
   +Triton(Ours) 10.93 29.30 41.35 66.62 46.49 43.91
   +Triton-C(Ours) 9.89 28.19 32.48 69.34 39.23 43.65
[…]
Triton-X(Ours) 20.85 40.77 39.12 73.00 54.87 43.98
Table 3: The experimental result on the HybriDialogue dataset with different baselines. The best results within each model are displayed in bold.
4.5 Experimental Setup
We illustrate the details of the experimental setup in Appendix D.
4.6 Experimental Result and Analysis
Tables 2 and 3 show the experimental results of different methods on two public datasets. We analyze the generated responses from the perspectives of text quality and factual consistency.
Text Quality
With the KG, the BLEU-4 and ROUGE-L scores fluctuate across two datasets. Our proposed Triton, the knowledge retriever, generally perform better than other knowledge-enhanced baselines, while we see a drop with Triton-C, the combination of dialogue rewrite and knowledge retriever. Our proposed framework, Triton-X, which consists of three components, achieves the highest BLEU and ROUGE-L scores, surpassing the SOTA G-retriever and demonstrating the superiority of our method.
There is a slight growth in PPL of smaller language models like ChatGLM and Flan-T5-Large: incorporating the KG decreases text fluency. After analysis, we found that the LLMs sometimes can not understand the KG well and thus generate unpolished responses. We also utilise Case 1 from Section 4.9 to illustrate the factors contributing to low PPL.
Overall, knowledge-augmented methods show variable performance across different LLMs in terms of BLEU-4 and ROUGE-L scores, and they tend to increase PPL slightly for smaller language models. Our proposed Triton-X gains the best performance in BLEU-4 on both datasets and ROUGE-L on HybriDialogue, indicating a higher quality in generating responses.
Factuality
With retrieved methods, the fact score rises in all LLMs, and Triton has a higher increase than KAPING and BM25. From Triton-C, we can see a continuous improvement in fact score, which means the rewritten dialogue also contributes a slight improvement to the fact score.
The NEIP is the proportion of atomic facts that can not be directly found in external knowledge to support or which are not a factual claim, and the lower one indicates the responses contain more verifiable facts. Compared to baseline LLMs, knowledge-augmented methods generally help reduce this proportion and produce more verifiable facts. From the tables, we observe a consistent decrease in NEIP for the Triton compared to KAPING and BM25. Furthermore, we noticed a significant improvement in NEIP with the rewritten dialogue. This suggests that rewriting the dialogue can substantially enhance factuality.
It is worth noting that the F1 score is not always consistent with the fact score. For instance, the trend of the F1 score shown in Flan-T5-XXL is almost opposite to the fact score on the OpendialKG dataset. One reason is that the reference responses lack explicit answers. Another reason is that the F1 score ignores the semantic aspects, and two sentences with the same entities can express contrary meanings.
In short, knowledge-augmented methods improve the factuality of LLMs, including the fact score and NEIP. Our proposed Triton series methods perform better than G-Retriever, KAPING and BM25 in the above metrics.
Model Scale
We conduct experiments with different LLMs. The Flan-T5 series LLMs (Large and XXL) are beneficial for us in analysing the correlation between size and performance.
As the model scale grows, fact score and NEIP exhibit differing trends across the two datasets—factuality declines on HybriDialogue, whereas the OpendialKG dataset shows the opposite pattern. The factuality is not always consistent with the model size, which contradicts our intuition.
HybriDialogue OpendialKG
[…]
w/o Retriever -1.1 -0.11 -1.1 +1.2
w/o Rewrite -2.3 +0.48 -1.46 +0.44
w/o Graph -3.41 -0.03 -1.00 -1.52
Table 4: Ablation study of Triton-X, and w/o means removing corresponding components.
4.7 Ablation Study
We conduct an ablation study with our proposed framework, shown in Table 4. We remove our components one by one from Triton-X to observe the performance. We only report factually related metrics as they are our aim. The decline in performance after removing either the rewrite or the retriever component highlights their importance.
4.8 Human Evaluation Results
HybriDialogue OpendialKG
Methods Cohe. Flu. Info. Cohe. Flu. Info.
ChatGLM-6B 1.90 1.99 1.71 1.99 1.91 1.11
+bm25 1.85 1.97 1.75 1.99 1.89 1.09
+SenEmb 1.84 1.98 1.71 1.98 1.95 1.18
+Triton 1.93 1.98 1.76 2.00 1.94 1.11
+Triton-C 1.94 1.96 1.85 1.98 1.97 1.28
Flan-T5-XXL 1.62 1.74 1.74 1.56 2.00 1.36
+bm25 1.67 1.72 1.80 1.38 1.94 1.49
+SenEmb 1.59 1.80 1.77 1.46 1.99 1.53
+Triton 1.59 1.81 1.77 1.52 1.98 1.39
+Triton-C 1.51 1.83 1.76 1.47 1.97 1.65
G-retriever 1.92 1.82 1.84 1.96 1.96 0.71
TritonX 1.96 1.81 1.82 2.00 2.00 0.60
Table 5: Human evaluation results for coherence (Cohe.), fluency (Flu.), and informativeness (Info.). The best results are marked within each category.
We report the human evaluation with the perspective: coherence, fluency and informativeness in Table 5. Each aspect is rated as one of three levels: positive, neutral, or negative (e.g., coherent, neutral, incoherent).
Our proposed Triton series methods demonstrate notable improvements in coherence across various settings, with the exception of the Flan-T5-XXL model. In terms of fluency, our methods achieve the highest scores for Flan-T5-XXL on HybriDialogue, as well as for ChatGLM and graph-based methods on OpendialKG. Overall, the proposed methods maintain strong performance in fluency and coherence. Additionally, the Triton series generally enhances informativeness.
4.9 Case Study
We use two cases that are beneficial for us in understanding our methods’ limitations. See more descriptions in Appendix H.
5 Conclusion
Given the limitations of previous methods in accurately retrieving knowledge triples and their challenges in handling dialogue tasks due to coreference issues, we propose a novel framework. This framework features a carefully designed knowledge retriever to identify valuable knowledge triples, an advanced dialogue rewriter to resolve coreference effectively, and a knowledge-enhanced dialogue generation component to improve factuality.
Building on this, we also adapt the fact score to make it suitable for evaluating the factuality of the dialogue response and verifying its validity in the dialogue task.
The results demonstrate that our method significantly enhances the factual accuracy of dialogue responses compared to other baselines, including the SOTA G-retriever.
We analyse some cases and see some limitations in our current work. We plan to explore the mechanism of LLM reactions to triples in the future.
Limitations
Although we propose a well-designed framework for dialogue-related tasks, some limitations remain.
Our current knowledge retrieval method relies on the exact entity overlap between the query and Wikidata. This approach lacks semantic understanding, making it inefficient in retrieving semantically related knowledge triples. Constructing a graph database from Wikidata could be an effective way to address this problem, but building such a large-scale database presents significant challenges.
Additionally, our approach relies on knowledge triples. However, the validity of certain triples, such as (“Montevideo”, “population”, “309331”) shown in the key knowledge of Case 1 in Section 4.9, will change over time. Retrieving such triples for LLMs could lead to outdated responses.
Ethical Statement
In this work, we follow the ethical standards in data use and research. The datasets used in this study, OpendialKG and HybriDialogue, are publicly available and have been widely used in research. We have not added any extra personal information to these publicly available datasets.
The data from human annotators was limited to some specific labels, and the results reported in this paper are based on statistical analysis. No personal or sensitive information was accessed during the research process.
Appendix A Prompts
The prompts of the knowledge-augmented approach are shown in Table 6. The factuality-related prompts are described in Table 8.
Dialogue Rewrite Prompt
You are tasked with resolving all pronouns and references in the given dialogue to their explicit entities. Use CoT (Chain of Thought) reasoning to identify what each pronoun or reference corresponds to. Do not answer any questions; your only goal is to perform co-reference resolution. Instructions: 1. Analyze the dialogue and process each turn in the conversation. 2. For every pronoun, ambiguous term, or reference, trace back in the conversation to determine its explicit entity or subject. 3. Clearly document your CoT reasoning for each resolution. 4. Provide the explicit reference for each pronoun or ambiguous term. Output Format: **Chain of Thought**: [Your reasoning process for resolving the references] **Resolved Dialogue**: [The dialogue with all pronouns and references resolved] Dialogue: {Dialogue}
Table 6: Prompt for rewriting the dialogue and generating a response with dialogue and knowledge triples.
Dialogue Response Generation Prompt Atomic Fact Splitting Prompt
Knowledge: {Selected Triples} Dialogue: {Dialogue Context} Given the above knowledge and dialogue, please respond to the input below and ensure the response is fluent and fact-consistent in English. Input: {Query} Response: … {Examples..} If the following input is an incomplete sentence or a phrase, please output it exactly as it is. Otherwise, if it is a complete sentence, split it into atomic sentences based only on the given information, without adding any additional information or making inferences: Input: {Response} Output …:
Table 7: The prompts for getting atomic facts and fact score.
Fact Score Prompt
Instruction: The statement is part of a response in a dialogue. Evaluate the statement strictly based on the provided knowledge source and dialogue history only. If the statement is not a factual claim (e.g., opinion, question, or unclear assertion), output: "no enough information." If it is a factual claim: Output true if the statement is directly supported by evidence in the knowledge source or dialogue history. Output false if the statement is directly contradicted by the knowledge source or dialogue history. Output no enough information if there is no direct evidence for or against the statement. Important: Do not use your intern knowledge or make inferences. Please only output your final answer and do not output any explanations. Evidence: {Wikipedia Passages} Dialogue history: {Dialogue} Speaker A: {Speaker} Statement: {Atomic Fact}
Table 8: The prompts for getting atomic facts and fact score.
Appendix B Datasets Details
This section provides a detailed description of the datasets, listed as follows:
OpenDialKG
It is an English knowledge-driven dialogue dataset that consists of a recommendation task related to movies and books and a chit-chat task related to sports and music. The dataset contains 13,802 samples. After processing, some samples are not valid in the task, so we remove them. Finally, we randomly selected 15% (1,962 samples) for the validation set, another 15% (1,973 samples) for the test set, and the remaining 70% (9,120 samples) for the training set. The dataset provides a number of triples extracted from Freebase Bast et al. (2014). These triples were collected several years ago. To address any potential issues with outdated information, we re-extract the triples from Wikidata. OpendialKG is released under CC-BY-NC-4.0 license [^2] https://github.com/facebookresearch/opendialkg, which permits non-commercial research use.
HybriDialogue
It is an open-domain and information-seeking dialogue dataset in English. It is constructed by splitting complex questions into multi-turn dialogue. The original dataset does not offer triples, so we collect them by matching the entities with triples from Wikidata. The training set contains 4,359 samples, while the validation set includes 242 samples, and the test set consists of 243 samples. HybriDial follows the MIT license Nakamura et al. (2022), which allows both commercial and non-commercial use.
OpendialKG Example
Query: Could you recommend some movies staring Chiaki Kuriyama?
Retrieved triples from BM25: (Chiaki Kuriyama, given name, Chiaki) (Chiaki Kuriyama, family name, Kuriyama) (Chiaki Kuriyama, place of birth, Tsuchiura) (Chiaki Kuriyama, occupation, film actor) (Chiaki Kuriyama, instrument, voice) …
Retrieved triples from KAPING: (Chiaki Kuriyama, given name, Chiaki) (Chiaki Kuriyama, occupation, model) (Chiaki Kuriyama, occupation, singer) (Chiaki Kuriyama, occupation, actor) (Chiaki Kuriyama, occupation, fashion model) …
Retrieved triples from Triton: (Kill Bill Volume 1, cast member, Chiaki Kuriyama), (Into the Sun, cast member, Chiaki Kuriyama), (Kagen no Tsuki, cast member, Chiaki Kuriyama), (Kamogawa Horumo, cast member, Chiaki Kuriyama), (Gonin, cast member, Chiaki Kuriyama) …
HybriDialogue Example
Query: Hi. Can you tell me who Judd Trump is?
Retrieved triples from BM25: (Judd Trump, given name, Judd)(Judd Trump, family name, Trump)(2021 Champion of Champions, winner, Judd Trump)(Judd Trump, victory, 2016 China Open)(Judd Trump, victory, 2011 China Open) …
Retrieved triples from KAPING: (Judd Trump, occupation, snooker player) (Judd Trump, nickname, The Ace) (Judd Trump, given name, Judd) (Judd Trump, family name, Trump) (Judd Trump, place of birth, Whitchurch) …
Retrieved triples from Triton: (2019 World Snooker Championship, winner, Judd Trump), (Judd Trump, award received, Snooker Hall of Fame), (Judd Trump, country of citizenship, United Kingdom), (Judd Trump, sport, snooker), (Judd Trump, award received, player of the year award) …
Table 9: The examples extracted from OpendialKG and HybriDialogue dataset. The triples are retrieved from Wikidata and selected using different methods.
Table 9 presents the two examples of OpendialKG and HybridDial datasets with selected triples from different methods.
Appendix C Examples from GPT-4o
We employ GPT-4o to annotate query-triple pairs and resolve dialogue coreference from the training dataset. In the end, we collected 29K samples of query-triple pairs and 35.4K samples of coreference resolution with reasoning steps. The examples for fine-tuning are listed in Tables 10 and  11.
Query Triple Relevance
Which teams has the Colorado Avalanche played against in the playoffs? (Colorado Avalanche, sport, ice hockey) Irrelevant
Hi! Do you know anything about Lutheran churches? (Lutheranism, described by source, Brockhaus and Efron Encyclopedic Dictionary) Irrelevant
Tell me more about Upper Canada District School Board? (Upper Canada District School Board, headquarters location, Brockville) Relevant
Hi, do you by any chance know anything about Vladimir Horowitz? (Stefanos Korkolis, student of, Vladimir Horowitz) Relevant
Table 10: The query and triple are sourced from the training dataset and Wikidata. The relevance is annotated by GPT-4o.
Section Content
Original Dialogue Speaker A: I like Lev Grossman books can you recommend some of his book Speaker B: Sure! He wrote The Magician’s Land and The Magicians. Speaker A: Oh, I think I’ve read The Magicians it a High Fantasy book.
CoT Speaker A mentions "Lev Grossman" as the author they like. This establishes that "his" in their request for book recommendations refers to Lev Grossman. Speaker B mentions "He wrote The Magician’s Land and The Magicians." The pronoun "He" refers to Lev Grossman, as he is the author being discussed. Speaker A uses "it" in "I think I’ve read The Magicians it a High Fantasy book." The pronoun "it" refers to "The Magicians," as it is the last explicitly mentioned book in the conversation.
Resolved Dialogue Speaker A: I like Lev Grossman books can you recommend some of Lev Grossman’s books Speaker B: Sure! Lev Grossman wrote The Magician’s Land and The Magicians. Speaker A: Oh, I think I’ve read The Magicians. The Magicians is a High Fantasy book.
Table 11: An example generated by GPT-4o illustrates how CoT reasoning works in coreference resolution.
Appendix D Experimental Setup
We provide detailed information regarding the experimental setup of our framework and the evaluation metrics used.
For fine-tuning the Triton, we only collect the query-triple pairs from the HybriDial dataset. We fine-tuned 3 epochs for either the LoRA method or the full-tuning method. The rank and alpha parameters for LoRA are both set to 8.
We utilise the multi-qa-mpnet-base-dot-v1 version for the baseline KAPING. As G-retriever is implemented using Llama2 7B Touvron et al. (2023), we adopt the same model in our Triton-X under identical settings to ensure fairness. For knowledge-augmented methods, the maximum number of selected triples is set to 5. Additionally, all experiments were run a single time.
Regarding the evaluation metrics, we utilised the evaluate package  [^3] https://huggingface.co/docs/evaluate to compute BLEU-4 and ROUGE-L scores. Besides, the PPL is calculated using GPT-2 Radford et al. (2019).
We fine-tune our relevance classifier using a single A100 GPU and our dialogue generation model with three A100 GPUs. We utilized two A100 GPUs for inference tasks. All A100 GPUs we used in this work have 80GB of memory. In total, our GPU runtime amounted to approximately 575 hours.
Appendix E Conventional Evaluation Metrics in Detail
We list all the details of conventional evaluation metrics as follows:
  • BLEU Reiter (2018), used in the machine translation initially, calculates the n-gram overlap between generated text and references and reflects the correlation between generated text and human writing.
  • ROUGE-L Lin (2004) is calculated based on the length of the longest common subsequence, which captures the word order and sentence-level structure from the ground truth.
  • PPL Jelinek et al. (1977) is widely used in measuring text fluency of generated output, in which calculation is based on the language model.
  • F1 Score Nan et al. (2021) is based on the precision and recall of the extracted entities matching between the generated response and ground truth.
Case 1 Case 2
Context: A: What is the capital and largest city in Uruguay? B: Montevideo is the capital and largest city in Uruguay. Query: What was the population of Montevideo, Uruguay, in 2011? Response: The population of Montevideo, Uruguay, in 2011 was 59,027. Key Knowledge: (Montevideo, population, 309331) New Response: Montevideo, population, 309331 Dialogue: A: What kind of game was “Team Fortress 2”? B: It’s a multiplayer first-person shooter. A: And what was it’s release date? Response: It was released on October 26, 2007. Rewritten Dialogue: B: What kind of game was “Team Fortress 2”? A: “Team Fortress 2” is a multiplayer first-person shooter. A: And what was “Team Fortress 2”’s release date? New Response: “Team Fortress 2” was released on November 20, 2007.
Table 12: The responses of cases were both generated by Flan-T5-Large. The new response was created using key knowledge or a rewritten dialogue. Triton selected the key knowledge.
Appendix F Details of Manual Annotation
We invited some students from Asian countries to annotate samples. All of them are well-educated and possess good English proficiency. The annotators were told that the annotated data would be used for research, and we got consent for data use.
To assess the accuracy of query-triple pair annotated from GPT-4o, we invite tree annotators to this task. We randomly selected 100 samples and instructed participants to determine whether the query and triple were relevant. The output is either relevant or irrelevant. In the beginning, two annotators have 0.83 in agreement and 0.66 in Cohen’s Kappa. We introduced another annotator to mediate the disagreement. We utilized the final annotated result to assess the accuracy of query-triple pairs, which is 0.81 in agreement and 0.62 in Cohen’s Kappa. These results indicate a high level of accuracy.
Three annotators were involved in the assessment of the fact score task. We used the fact score prompt, described in Table 8, as an instruction. Two annotators independently assessed the fact score for 100 samples from the HybriDial and OpendialKG datasets respectively. Initially, the raw agreement scores were 0.76 for HybriDial and 0.81 for OpenDialKG. Their inter-annotator agreement, measured by Cohen’s Kappa, was 0.613 for OpendialKG and 0.705 for HybriDial. To resolve discrepancies, we introduced a third annotator to mediate disagreements. Finally, we report the final agreement and Cohen’s Kappa score with the evaluation model, as shown in Table 1.
Two annotators were involved in the dialogue human evaluation. The accuracy of human annotations is consistently high across all aspects, with HybriDial achieving 0.81 for coherence, 0.88 for fluency, and 0.89 for informativeness, while OpendialKG reaches 0.83, 0.96, and 0.80 respectively, indicating the overall reliability of manual evaluations. We finally report the average score in our dialogue human evaluation.
Appendix G Baseline Descriptions
LLMs used in this work are presented as follows:
  • ChatGLM Zeng et al. (2022) is an open-source and bilingual language model, based on GLM Du et al. (2021) architecture. The technique of ChatGLM-6B is similar to ChatGPT, optimized in dialogue tasks.
  • Flan-T5 Chung et al. (2022) proposed several instruction-based LLMs, scaling the number of tasks and model size and fine-tuning in the chain-of-thought data. We adopt Flan-T5-XXl (11B), Flan-T5-Large (780M) and Flan-T5-Small (80M) as baselines in our work.
The baseline methods are described as follows:
  • No Knowledge: We feed the prompt into LLMs to generate dialogue responses without external knowledge.
  • BM25 Robertson et al. (2009): It is the ranking function between a query and several documents widely used in information retrieval. In our experiments, we also apply RefinED to retrieve triples for BM25 and replace documents with triples for retrieval.
  • KAPING Baek et al. (2023): It is a popular knowledge-augmented method for question answering. Since the authors have not released the official code, we follow the experimental setup of the KAPING method: we first extract entities from the query using REfinED, rank the query with triples using MPNet, select the top-N most relevant triples, and supplement them as prompts to generate dialogue responses.
  • G-Retriever He et al. (2024): It is a retrieval-augmented generation method for textual graphs that selects informative subgraphs via a Prize-Collecting Steiner Tree, encodes them with GNNs, and generates responses with high accuracy, outperforming strong baselines and reducing hallucinations.
  • Triton(Ours): We use GPT-4o to generate high-quality query-triple pairs based on the training dataset and fine-tune Llama3 8B base model with collected samples. The model acts as a triple retriever to help select valuable triples for generating dialogue responses.
  • Triton-C(Ours): We fine-tune a coreference resolution model based on Llama 8B base model using reasoning distillation. Our approach integrates dialogue rewriting with entity linking for enhanced performance.
  • Triton-X(Ours): It is the framework that consists of our well-designed knowledge retriever, dialogue rewrite and graph-based module.
Appendix H Details of the Case Study
The examples of case studies are shown in Table 12.
The topic in Case 1 is about the population of Montevideo, and our methods do help select the relevant triples. However, the LLM then generates one of triples directly, which is not fluent and has a higher PPL i.e. lower quality.
Case 2 is a dialogue that asks about the release date of “Team Fortress 2”, and the original response responds to the dialogue and is close to the writing of reference. After rewriting the dialogue, the LLM generates a response that differs from the reference, causing a lower BLEU-4 score.

### v2

@@
-Improving Factuality for Dialogue Response Generation via Graph-Based Knowledge Augmentation
+Improving Factuality for Dialogue Response Generation via Graph-Based Augmentation
 Abstract
 Large Language Models (LLMs) succeed in many natural language processing tasks.
-However, their tendency to hallucinate — generate plausible but inconsistent or factually incorrect text — can cause problems in certain tasks, including response generation in dialogue.
-To mitigate this issue, knowledge-augmented methods have shown promise in reducing hallucinations.
-Here, we introduce a novel framework designed to enhance the factuality of dialogue response generation, as well as an approach to evaluate dialogue factual accuracy.
-Our framework combines a knowledge triple retriever, a dialogue rewrite, and knowledge-enhanced response generation to produce more accurate and grounded dialogue responses.
-To further evaluate generated responses, we propose a revised fact score that addresses the limitations of existing fact-score methods in dialogue settings, providing a more reliable assessment of factual consistency.
+However, their tendency to hallucinate — generate plausible but inconsistent or factually incorrect text — can cause significant problems in certain tasks, including response generation in dialogue.
+To mitigate this issue, we propose two novel graph knowledge-augmented frameworks, Dialogue Response Generation via Textualised Graphs (TG-DRG) and Graph-Aware Dialogue Response Generation (GA-DRG), which combine reasoning-guided dialogue reformulation, dialogue sense knowledge selection, and graph-enhanced response generation to improve the factuality of dialogue responses.
+To evaluate the factuality of generated responses, we propose a dialogue fact score that addresses the limitations of existing fact-score methods in dialogue settings, providing a more reliable assessment of factual consistency.
 We evaluate our methods using different baselines on the OpendialKG and HybriDialogue datasets.
-Our methods significantly improve factuality compared to other graph knowledge-augmentation baselines, including the state-of-the-art G-retriever.
-The code will be released on GitHub  [^1] https://github.com/XiangyanChen/Triton.
-Improving Factuality for Dialogue Response Generation via Graph-Based Knowledge Augmentation
-Xiangyan Chen1  Yujian Gan1  Matthew Purver1,2 1Queen Mary University of London, UK 2Institut Jožef Stefan, Slovenia {xiangyan.chen, y.gan, m.purver}@qmul.ac.uk
+Our methods noticeably improve factuality compared to other graph knowledge-augmentation baselines, including the state-of-the-art G-retriever, achieving improvements of 3.47% on OpendialKG and 3.12% on HybriDialogue in terms of dialogue fact score.
+The code will be released on GitHub [^1] https://github.com/XiangyanChen/DRG.
 1 Introduction
-Figure 1: Examples comparing an LLM’s dialogue response generation when utilising retrieved triples and a rewritten dialogue versus generating responses without external knowledge or dialogue rewriting.
-The knowledge triple retriever module assists in selecting the triples most relevant to the query.
 Large Language Models (LLMs) have been shown to perform powerfully on Natural Language Processing (NLP) tasks.
 Despite their general superiority, LLMs will generate some plausible but fact-inconsistent text, namely hallucination.
-Flawed pre-training data, model bias and randomness in inference are the factors contributing to hallucinations Zhang et al.
-(2023); Huang et al.
-(2023).
+Flawed pre-training data, model bias and randomness in inference are the factors contributing to hallucinations (Zhang et al.
+2023; Huang et al.
+2023).
 In dialogue response generation, generating incorrect responses will mislead people and have a negative impact on society.
-A number of methods have been proposed to enhance the factuality of language models.
-Among them, knowledge-augmented inference methods aim to do this by adding knowledge triples into the prompt.
-This has been effective in tasks such as question-answering Baek et al.
-(2023); Sen et al.
-(2023); Wu et al.
-(2023).
-However, these works Baek et al.
-(2023); Wu et al.
-(2023) tend to use simple sentence embedders to encode knowledge and determine the match between a query and knowledge based on similarity, rather than employing a specialised model designed to assess the relevance between the query and knowledge.
+A number of approaches have been proposed to enhance the factuality of LLMs.
+Among them, graph knowledge-augmented methods aim to improve factual accuracy by incorporating knowledge graphs.
+This has been effective in tasks such as Question Answering (QA) (Baek, Aji, and Saffari 2023; Wu et al.
+2023; He et al.
+2024).
+Figure 1: An example illustrates how our frameworks TG-DRG and GA-DRG, and dialogue fact score work.
+Recently, Baek, Aji, and Saffari (2023) utilised sentence embedders to encode knowledge and determine the match between a query and knowledge based on similarity.
 He et al.
-(2024) proposed a graph-based method called G-retriever, which employs the Prize-Collecting Steiner Tree to help select knowledge triples.
+(2024) proposed a graph-based method called G-retriever, which employs the Prize-Collecting Steiner Tree (PCST) to help construct a subgraph.
 However, applying these methods to dialogue response generation is challenging due to their lack of consideration for dialogue context.
-Dialogues frequently involve intricate coreference structures, which complicate entity linking and hinder the LLMs’ comprehension, ultimately compromising the factual consistency and quality of the generated responses.
-To address above limitations, we propose a novel framework to improve dialogue factuality that consists of three key components: a knowledge triple retriever, a dialogue rewriting and knowledge-enhanced dialogue response generation.
-The knowledge triple retriever is fine-tuned on our well-collected samples, which helps select the valuable triples for the dialogue.
-The dialogue rewrite is based on Chain of Thought (CoT), which resolves the dialogue coreference.
-The knowledge-enhanced dialogue generation is proposed to improve the dialogue factuality, based on two kinds of approaches: prompt-based and graph-based.
-Figure 1 illustrates how the knowledge-augmented generation works.
-The query asks about what movies star “Diane Lane”, and the initial response generated by the LLM contains factual inaccuracies.
-After applying a dialogue rewrite module that resolves coreference, a triple retriever helps select the most relevant N triples, including (“Nights in Rodanthe”, “cast member”, “Diane Lane”).
-With these corrected triples, the model generates an accurate response.
-To evaluate factuality for responses, we primarily rely on the fact score Min et al.
-(2023).
-However, the original design of the fact score does not consider dialogue context and situations where the knowledge source is unavailable, making it difficult to evaluate dialogue responses.
-To address this limitation, we adapt the fact score to the dialogue settings and assess model-human agreement.
-Annotation results show that our adapted fact score achieves substantial model-human agreement.
+Dialogues frequently involve intricate coreference structures, which hinder the LLMs’ comprehension, ultimately compromising the factual consistency and quality of the generated responses.
+In addition, evaluating factuality for dialogue responses is a challenge.
+F1-score relying on entities overlapping has limitations on open-domain tasks.
+More advanced FactScore (Min et al.
+2023) evaluates facuality by splitting long-text into smaller atomic facts, but it does not consider dialogue context and situations where the knowledge source is unavailable, making it difficult to evaluate dialogue responses.
+To overcome the limitation on generating factually correct responses, we propose two frameworks: (1) Dialogue Response Generation via Textualised Graphs (TG-DRG), where generation requires no training; and (2) Graph Aware Dialogue Response Generation (GA-DRG), which explicitly models node relationships, designed to improve the factual accuracy of dialogue generation.
+Both frameworks are built around reasoning-guided dialogue reformulation and dialogue sense knowledge selection.
+The latter helps select the most valuable knowledge based on the dialogue.
+We use an example shown in Figure 1 to illustrate how our frameworks work.
+The response to the dialogue is incorrect due to the pronouns, e.g.
+“her”.
+After applying Chain of Thought (CoT), the pronoun is solved, benefiting further knowledge-augmented dialogue response generation.
+To address limitations in evaluating factual accuracy, we propose the Dialogue Fact Score, as illustrated in Figure 1.
+In this case, the FactScore can not work without a dialogue context.
+Our metric accounts for both the dialogue context and scenarios in which the knowledge source is partially or entirely unavailable.
 Additionally, we propose Not Enough Information Proportion (NEIP) to evaluate dialogue factuality comprehensively.
 It is a metric that quantifies the proportion of atomic facts in a response that cannot be verified, such as opinions or hallucinated content that can not find any evidence to verify.
-We utilize the F1 score as a supplementary metric to assess factual accuracy and evaluate response quality through BLEU Reiter (2018), ROUGE-L Lin (2004), and Perplexity (PPL) Jelinek et al.
-(1977).
-Additionally, we perform human evaluations to further assess response quality.
-We compare our proposed method not only with standalone LLMs but also with knowledge-enhanced approaches such as KAPING, the BM25 algorithm, and the State-of-the-Art (SOTA) G-retriever He et al.
-(2024).
+We assess model-human agreement of dialogue fact score.
+Annotation results show that our dialogue fact score achieves substantial model-human agreement.
+We compare our proposed method not only with standalone LLMs but also with knowledge-enhanced approaches such as KAPING (Baek, Aji, and Saffari 2023), the BM25 algorithm, and the State-of-the-Art (SOTA) G-retriever (He et al.
+2024).
 Experimental results show that our framework consistently outperforms both existing knowledge-enhanced methods and the current best-performing G-retriever.
@@
   1. 1.
-    We propose a novel framework designed to enhance the factual accuracy of dialogue response generation.
-This framework incorporates three key components: a knowledge triple retriever that selects valuable triples, a dialogue rewrite that resolves coreference, and a knowledge-enhanced mechanism that effectively boosts overall factuality.
+    We propose the dialogue fact score to evaluate the factuality of dialogue systems automatically.
+The evaluation is fine-grained, enabling a more reliable assessment of factual consistency in dialogue systems.
   2. 2.
-    We adapt the fact score, originally designed for biography generation, to evaluate the factuality of dialogue systems, ensuring its validity within our task.
-The evaluation is fine-grained, enabling a more reliable assessment of factual consistency in dialogue systems.
+    We propose an unsupervised framework, TG-DRG, designed to enhance the factual accuracy of responses without generation training.
+This framework adopts a textualised graph to generate responses built around reasoning-guided dialogue reformulation and dialogue sense knowledge selection.
+We validate TG-DRG against various comparable baselines; the experimental results show that TG-DRG can improve factuality noticeably.
   3. 3.
-    We validate our methods against various baselines using two public dialogue datasets.
-Experimental results demonstrate that our approach significantly outperforms existing methods in factuality while keeping good quality at the generated response.
+    We propose a supervised framework, GA-DRG.
+Different from TG-DRG, GA-DRG utilises a graph-based neural network to capture the node relationships, making it more efficient in improving factuality when generating dialogue responses.
+The experimental results show that GA-DRG outperforms the SOTA G-retriever in dialogue fact score with an improvement of 3.47% on OpendialKG and 3.12% on HybriDialogue.
+Figure 2: The workflow of our proposed frameworks, TG-DRG and GA-DRG, begins with the dialogue.
+A CoT process first reformulates the dialogue.
+Based on the refined dialogue, the most valuable knowledge triples are then selected by the dialogue sense knowledge selection module.
+Starting from Step 4, the workflow diverges into two distinct paths for dialogue response generation: the blue solid arrows represent the TG-DRG path, while the purple dashed arrows indicate the GA-DRG path.
 2 Related Work
 2.1 LLM Hallucinations and Mitigation Methods
-LLMs are trending in NLP, whose architecture can mainly be classified as encoder-only Devlin (2018), decoder-only Brown (2020); Dubey et al.
-(2024), and encoder-decoder Chung et al.
-(2024).
-However, hallucination exists in these LLMs widely, which is the phenomenon that language models generate coherent but fact-inconsistent text.
-Many reasons contribute to hallucination.
-Huang et al.
-(2023) declared one potential reason is that the data in the pre-training of LLMs is incomplete, incorrect or outdated.
-Other reasons for hallucinations are the bias of LLMs and randomness in inference.
-Zhang et al.
-(2023) categorise hallucinations as input-conflicting, context-conflicting and fact-conflicting, with recent focus on the latter.
-Several knowledge-based methods have been proposed to mitigate fact-conflicting hallucinations Agrawal et al.
-(2023), which generally fall into three categories: knowledge-aware inference, knowledge-aware training, and knowledge-aware validation.
-For instance, Baek et al.
-(2023) introduced KAPING, a prompt-based framework for question answering that retrieves knowledge triples based on embedding similarity.
-Building upon retrieval techniques, Sen et al.
-(2023) combined Knowledge Graph (KG) retrieval with language model reasoning to enhance performance on complex questions.
-Similarly, CRAG Yan et al.
-(2024) employs an evaluator to assess generation quality and, if necessary, refines outputs via web search, showing significant improvements across various generation tasks.
-Another notable approach is Self-RAG Asai et al.
-(2023), which retrieves relevant knowledge, generates an initial answer, and then refines it through self-critique, achieving strong results in multiple question answering benchmarks.
+LLMs are trending in NLP. However, hallucination exists in these LLMs widely, which is the phenomenon that language models generate coherent but fact-inconsistent text.
+Several knowledge-based methods have been proposed to mitigate fact-conflicting hallucinations (Agrawal et al.
+2023), which generally fall into three categories: knowledge-aware inference, training, and validation.
+For instance, Baek, Aji, and Saffari (2023) introduced KAPING, a prompt-based framework for QA that retrieves knowledge triples based on embedding similarity.
+Building upon retrieval techniques, Sen, Mavadia, and Saffari (2023) combined Knowledge Graph (KG) retrieval with language model reasoning to enhance performance on complex questions.
+Similarly, CRAG (Yan et al.
+2024) employs an evaluator to assess generation quality and, if necessary, refines outputs via web search, showing significant improvements across various generation tasks.
+Another notable approach is Self-RAG (Asai et al.
+2023), which retrieves relevant knowledge, generates an initial answer, and then refines it through self-critique, achieving strong results in multiple question answering benchmarks.
 In a related development, He et al.
-(2024) proposed G-Retriever, which integrates a graph-based encoder with the Prize-Collecting Steiner Tree algorithm to encode retrieved graph knowledge effectively, standing out among graph-enhanced knowledge retrieval methods.
+(2024) proposed G-Retriever, which integrates a graph-based encoder with the PCST algorithm to encode retrieved KG effectively, standing out among graph-enhanced knowledge retrieval methods.
+However, applying QA-based approaches to multi-turn dialogue generation is limited by their lack of coreference modelling, which is essential for capturing discourse-level dependencies and ensuring factual consistency.
 2.2 Evaluation of Factuality
-Dialogue response generation differs from question answering in that it must not only take the dialogue context into account when generating responses, but also must be evaluated based on the factuality and quality of the responses, rather than just the accuracy of answers.
-Previous dialogue factuality evaluation mainly focuses on humans Ni et al.
-(2023); Li et al.
-(2022); Yu et al.
-(2022), and it is inefficient.
+Dialogue response generation not only needs to consider the dialogue context when generating responses but also needs to be evaluated based on the factuality and quality, rather than just the accuracy of answers of QA.
+Previous dialogue factuality evaluation mainly focuses on humans (Ni et al.
+2023; Li et al.
+2022; Yu et al.
+2022), which is inefficient.
 Automatic evaluating factuality is challenging, however; current methods are either reference-based or reference-free.
-Several question-answering datasets provide factual references in the form of entities; reference-based evaluation can then be based directly on entity matching.
-Other NLP datasets offer text references, so an alternative reference-based method involves matching the extracted entities between the generated text and these references Nan et al.
-(2021).
+Several QA datasets provide factual references in the form of entities; reference-based evaluation can then be based on entity matching.
+Other NLP datasets offer text references, so an alternative reference-based method involves matching the extracted entities between the generated text and these references (Nan et al.
+2021).
 Instead, dialogue datasets provide reference responses.
-To evaluate the factual accuracy of the generated text, we can therefore compare the entities extracted from the generated response with those in the reference response, measured as F1 score.
+To evaluate the factual accuracy of the generated text, we can therefore compare the entities extracted from the generated response with those in the reference response, measured as the F1 score.
 However, relying solely on this can be inadequate, as the reference may not always provide explicit answers.
 It is crucial to consider additional metrics to ensure a comprehensive evaluation.
-Reference-free methods can be categorised into uncertainty estimation Farquhar et al.
-(2024) and external knowledge-based approaches, with the former often limited by its reliance on the generation model’s own confidence.
-The fact score Min et al.
-(2023), which falls into the latter category, is a metric designed to measure fact consistency in long-form text.
+Reference-free methods can be categorised into uncertainty estimation (Farquhar et al.
+2024) and external knowledge-based approaches, with the former often limited by its reliance on the generation model’s own confidence.
+The FactScore (Min et al.
+2023), which falls into the latter category, is a metric designed to measure fact consistency in long-form text.
 The process begins with an LLM breaking down the text into fine-grained sentences called atomic facts.
@@
 The FactScore is calculated based on the precision.
-However, the fact score is not directly suitable for our work for two key reasons: First, it evaluates only the generated text without taking the dialogue history into account, which is crucial when assessing the factuality of dialogue responses.
-Second, it does not account for the possibility that generated responses may lack external knowledge sources for verification.
-Figure 2: The workflow of dialogue response generation and factuality evaluation.
-The left part of the figure is our framework.
-It starts with rewriting the dialogue, and then the knowledge triples will be retrieved from KG.
-The triple retriever helps select valuable triples.
-The colour-coded arrows marked as Serial Number 6 represent different knowledge-enhanced methods: the orange arrow indicates the graph-based approach, while the blue arrow corresponds to the prompt-based approach.
-The right part is factuality evaluation, in which the response will be broken down into individual and verifiable facts, checked using both the LLM and external knowledge sources.
+However, the FactScore is not directly suitable for our work for two reasons: First, it evaluates only the generated text without taking the dialogue history into account, which is crucial when assessing the dialogue factuality.
+Second, it does not account for the possibility that generated responses may lack external knowledge for verification.
 3 Methodology
-The diagram in Figure 2 illustrates the workflow of dialogue generation (left side) and factuality evaluation (right side).
-Our proposed framework consists of three key components: (1) A well-designed knowledge triple retriever and see Section 3.2 for details, (2) A CoT-based dialogue rewriter, described in Section 3.3, and (3) the knowledge-enhanced dialogue response generation (Section 3.2).
-For factuality evaluation (Section 3.5), we adopt the fact score as a primary metric.
-However, as the original fact score does not fully capture certain aspects of dialogue evaluation, we introduce modifications to better align with our objectives.
+To address the limitations of prior works (He et al.
+2024; Baek, Aji, and Saffari 2023), which often overlook complex pronoun resolution in dialogue settings and struggle to generate responses with high factual accuracy, we propose two novel frameworks: (1) Dialogue Response Generation via Textualised Graphs (TG-DRG) without training a generation model; and (2) Graph-Aware Dialogue Response Generation (GA-DRG), which leverages graph structures to better capture relationships between nodes.
+Both of them are built around a reasoning-guided coreference resolution and a dialogue sense knowledge selection.
+TG-DRG includes two variants based on the use of dialogue reformulation: TG-DRG-R incorporates a reformulation module, while TG-DRG-NR bypasses it.
+The overall workflow for our frameworks is illustrated in Figure 2.
+In addition, to automatically evaluate factuality in dialogue systems, we propose the Dialogue Fact Score, a metric designed to assess factual accuracy specifically within dialogue contexts, thereby addressing the limitations of existing evaluation methods for open-domain dialogue factuality.
 3.1 Task Formulation
-Our task is to generate a response $s$ given a dialogue context $c$, query $x$ and set of triples $f=\{(h_{1},r_{1},t_{1}),(h_{2},r_{2},t_{2}),...,(h_{n},r_{n},t_{n})\}$ from an encyclopedia-based KG, where $h$ and $t$ represent the head and tail entities, and $r$ denotes the relation between them.
-3.2 Knowledge Retriever
-The KG is a collection of triples, and these triples are the form of head entity, relation, and tail entity.
-In encyclopedia-based KGs, like Wikidata Vrandečić and Krötzsch (2014), a triple usually indicates a fact.
-We aim to retrieve triples from the KG for the query in dialogue and incorporate triples into prompts as a supplement.
-We first extract entities from the query using the entity linking approach, and then, we match all entities with the triples from the KG when the entity equals the head or tail entity.
-Since our entities are sourced from Wikidata, and ReFinED Ayoola et al.
-(2022) is an entity linking method based on corresponding Wikidata entities, we adopt it for our retrieval process.
-Knowledge retriever gathers all triples associated with the entities extracted from the query, but not all of these are relevant to the query.
-Irrelevant triples introduce textual noise, which has a negative impact on text generation.
-Thus, it is crucial to select the most relevant triples for the query.
-Traditional methods rank triples based on word frequency using BM25 Robertson et al.
-(2009), which primarily captures word overlap.
-Alternatively, sentence embedding models like MPNet Song et al.
-(2020), as employed by KAPING Baek et al.
-(2023), encode semantic information but in a relatively straightforward manner.
-To retrieve delicate knowledge, we collect query-triple pairs to fine-tune LLM as a knowledge matcher.
-Data Collection
-Manual data annotation is both time-consuming and labour-intensive.
-Given GPT’s strong performance in NLP tasks, we adopt it as an annotation tool.
-Our approach extracts query-triple pairs from the training dataset and prompts GPT-4o to determine their relevance.
-The model classifies each pair as either relevant or irrelevant.
-Based on our manual assessment, GPT-4o demonstrates high accuracy in determining query-triple relevance, described in Appendix F.
-Fine-Tuning
-After collecting these data, we fine-tune an LLM as the binary classifier.
-In our work, we select the Llama3 8B model Dubey et al.
-(2024).
-We combined LoRA Hu et al.
-(2021) with fine-tuning, which is an efficient fine-tuning method that adopts low-rank modules into transformer layers.
-Cross-entropy loss is used to optimize the model during fine-tuning, and it is formulated as follows:
-$\mathcal{L}=-\sum_{i=1}^{N}y_{i}\log(p_{i})$ (1)
-where $y_{i}$ denotes the true label, $p_{i}$ is defined as the predicted probability and $N$ is the number of label classifications.
-With the fine-tuned LLM $\mathcal{M}_{t}$, the query $x$ and the triple $f_{i}$, the probability of relevance can be defined as:
-$P(y=\text{relevant}|x,f_{i})=\mathcal{M}_{\text{t}}(x,f_{i})$ (2)
-Then, the selected triples are formulated as:
-$f_{\text{sorted}}=\text{sort}_{\text{desc}}(P,f,x)$ (3)
-where $\text{sort}_{\text{desc}}$ denotes the sort function descending by $P$, and $P$ is the relevance probability distribution.
-3.3 Dialogue Rewrite
-Unlike in question answering tasks, dialogue response generation involves coreferences that pose challenges for both entity linking and response generation.
-An intuitive approach to improve entity linking is to expand the scope for extracting entities from the dialogue context.
-However, this approach may also introduce irrelevant entities, negatively impacting performance.
-To resolve coreference and improve the performance of entity linking and dialogue response generation, we propose a CoT-based coreference resolution.
-Data Collection
-We prompt GPT-4o with dialogue to resolve the conversation in a zero-shot CoT setting, as described by Kojima et al.
-(2022).
-Step-by-step reasoning in LLMs enhances the generation of accurate results and aids in fine-tuning smaller language models.
-Fine-Tuning
-We fine-tune the Llama3 8B model using LoRA on collected coreference resolution samples.
-The reason for selecting LoRA for this task is that it will not destroy the LLMs reasoning capability obtained from pre-training, making it perform better on this task.
-Cross-entropy loss is also selected for this task.
-After fine-tuning, we prompt LLM to resolve coreference, and the dialogue is rewritten. The rewrite and corresponding sorted triples are formulated as follows:
-$c^{\prime},x^{\prime}=\mathcal{M}_{\text{f}}(p_{\text{rewrite}},c,x)$ (4)
-$f_{\text{sorted}}^{{}^{\prime}}=\text{sort}_{\text{desc}}(P,f,x^{\prime})$ (5)
-where $p_{\text{rewrite}}$, described in Table 6 of Appendix A, is the prompt for rewriting dialogue. $\mathcal{M}_{\text{f}}$ is the LLM for dialogue rewriting.
-3.4 Knowledge-Based Dialogue Response Generation
-To mitigate hallucinations of LLM and have a factual response, we offer two patterns of generating dialogue responses given selected knowledge triples: prompt-based and graph-based methods.
-Prompt Based Generation
-Given dialogue context $c$, query $x$, rewritten dialogue $c^{\prime}$ and $x^{\prime}$, as well as selected triples $f_{sorted}$ and $f_{sorted}^{{}^{\prime}}$, the generation of knowledge-augmented dialogue responses and their counterparts with the rewritten dialogue are formulated as follows:
-$s=\mathcal{M}_{\text{g}}(p_{\text{gen}},c,x,f_{\text{sorted}})$ (6)
-$s^{{}^{\prime}}=\mathcal{M}_{\text{g}}(p_{\text{gen}},c^{\prime},x^{\prime},f_{\text{sorted}}^{{}^{\prime}})$ (7)
-where $p_{\text{gen}}$ denotes the prompt for generating the response. $\mathcal{M}_{\text{g}}$ is the LLM for response generation. Table 6 in Appendix A describes the details of the prompt template.
-Graph-Based Generation
-Prompt-based generation can effectively improve factuality with training, while it is unable to capture the connectivity information within a Knowledge Graph.
-To address this limitation, we encode the knowledge graph with GNN (Graph Netural Network).
-We first convert the sorted triples $f^{{}^{\prime}}_{sorted}$ into a graph $\mathcal{G}$.
-Then the graph and dialogue encoding are defined as follows:
-$v=E_{graph}(\mathcal{G})$ (8)
-$z=\mathcal{M}(c^{\prime},x^{\prime})$ (9)
-where $E_{graph}$ is GNN. $\mathcal{M}$ represents an arbitrary LLM. $v$ and $z$ denote the graph and text embedding, respectively.
-We generate the dialogue response by decoding the concatenation of the graph and text embedding, formulated as follows:
-$s_{g}=Decoder([v;z])$ (10)
-where ; denotes concatenation and $s_{g}$ is the response generated by graph-based method. To preserve the capability of the LLM, we combine LoRA with cross-entropy to train our graph model.
-3.5 Evaluation Metrics on Fact
-Fact score Min et al.
-(2023) measures the fact consistency for long-form text.
-As we mentioned in Section 2.2, it does not consider the dialogue history and the possibility that the generated response may lack the knowledge source to support it.
-The atomic facts and fact verification prompts of fact score are modified in this paper.
-In the revised atomic fact-splitting prompt, only complete sentences are permitted to be split, and the model is prohibited from adding any additional information to the atomic facts.
-We incorporate dialogue history and introduce the “Not Enough Information” label for the revised fact verification prompt.
-If an atomic fact is not a factual claim or lacks direct support from any known sources, the model will output ‘Not Enough Information”.
-The detailed prompts for atomic fact-splitting and fact verification are provided in Table 8 of Appendix A.
-Since the original fact score framework does not specify how to get knowledge sources for dialogue, we extract entities from atomic facts and query as Wikipedia titles to search related passages as knowledge sources.
+We define the conversation history as a sequence of utterances $\mathcal{U}=\{\text{U}_{1},\text{U}_{2},\dots,\text{U}_{t-1}\}$, where $\text{U}_{i}$ denotes the utterance at the $i^{th}$ turn, either from the user or the system.
+In our setup, external knowledge is modeled as a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}=\{v_{1},v_{2},\dots,v_{n}\}$ represents the set of nodes, which are entities, and $\mathcal{E}=\{e_{1},e_{2},\dots,e_{m}\}$ represents the set of edges, which are relations.
+The graph is constructed from an encyclopedia-based knowledge source, with each knowledge fact represented as a triple $(s,p,o)$, where $s,o\in\mathcal{V}$ denote the head and tail nodes, respectively, and $p\in\mathcal{E}$ denotes the edge (relation) connecting them.
+Our task is to generate a factually correct response $\text{U}_{t}$ conditioned on the preceding utterances $\mathcal{U}$ and the knowledge graph $\mathcal{G}$.
+3.2 Reasoning-Guided Dialogue and Knowledge Integration
+We introduce a CoT-driven approach to dialogue reformulation, which systematically resolves coreferences step-by-step to produce high-quality dialogues.
+Additionally, we propose a dialogue sense knowledge selection method that constructs dialogue-relevant graphs to enhance factual accuracy.
+The reformulated dialogue enhances the effectiveness of knowledge selection and further mitigates hallucinations.
+Reasoning-Guided Dialogue Reformulation
+Zero-shot coreference resolution often struggles with resolving complex pronoun references in conversations due to the intricate reasoning involved.
+To address this challenge, we leverage the reasoning capabilities of LLMs to perform step-by-step resolution for each turn of the dialogue.
+This iterative approach not only simplifies the understanding of complex coreference relationships but also enhances the model’s overall coreference resolution ability.
+Specifically, we employ the CoT technique to guide the model through incremental reasoning, resulting in more accurate dialogue reformulation and improved factual consistency.
+$\mathcal{U^{{}^{\prime}}}=f^{\text{ref}}_{\theta}(p_{\text{reform}},\mathcal{U}),$ (1)
+where $p_{\text{reform}}$ is the prompt guiding the dialogue reformulation process, and $f^{\text{ref}}_{\theta}$ denotes an LLM that transforms the original dialogue $\mathcal{U}$ into its reformulated form $\mathcal{U}^{\prime}$.
+Dialogue Sense Knowledge Selection
+The external knowledge graph $\mathcal{G}$ contains noise and dialogue-irrelevant triples, making the LLM difficult to focus on the useful nodes when generating responses.
+To improve the quality, we propose a dialogue sense selection method relying on an LLM, which helps select the most dialogue-relevant triples and form a dialogue sense graph.
+Specifically, we firstly encode the dialogue $\mathcal{U}$ and the triple $(s,p,o)$, respectively, into vector representations by the encoder $\psi$ and $\phi$.
+The function $f^{\text{sel}}_{\theta}$ denotes the LLM, which computes the score based on these representations.
+The formulation of selection probability is listed as follows:
+$P(y=\text{rel}\mid\mathcal{U^{\prime}},(s,p,o))=\sigma\left(f^{\text{sel}}_{\theta}\left(\psi(\mathcal{U}^{(\cdot)}),\phi(s,p,o)\right)\right),$ (2)
+where $\mathcal{U}^{(\cdot)}\in\{\mathcal{U},\mathcal{U}^{{}^{\prime}}\}$, rel denotes relevance. $\sigma$ is the activation function that maps the output to the probability.
+We select the N triples based on the probability from the knowledge graph $\mathcal{G}$, and the process is defined as follows:
+$\mathcal{G}_{\mathcal{U}^{(\cdot)}}=\text{Top}_{N}\left(\mathcal{E},\ \left\{P\left(\text{rel}\mid\mathcal{U}^{(\cdot)},\ (s,p,o)\right)\right\}_{(s,p,o)\in\mathcal{E}}\right),$ (3)
+where $\mathcal{G}_{\mathcal{U}^{(\cdot)}}\in\{\mathcal{G}_{\mathcal{U}},\mathcal{G}_{\mathcal{U}^{{}^{\prime}}}\}$ is the dialogue sense graph constructed from the selected triples.
+We select the LLaMA 3 8B model for dialogue reformulation and dialogue sense knowledge selection to achieve a balance between capability and efficiency.
+To enhance performance and leverage the strengths of advanced models, we fine-tune the LLaMA models using generated samples from GPT-4o, employing a combination of cross-entropy loss and Low-Rank Adaptation (LoRA) techniques (Hu et al.
+2021).
+3.3 Graph Knowledge-Driven Response Generation
+The dialogue sense graph facilitates generating factual responses. We propose two graph-based generation methods.
+Dialogue Response Generation via Textualised Graphs
+Given the dialogue $\mathcal{U}^{(\cdot)}\in\{\mathcal{U},\mathcal{U}^{{}^{\prime}}\}$, as well as the dialogue sense graph $\mathcal{G}_{\mathcal{U}^{(\cdot)}}\in\{\mathcal{G}_{\mathcal{U}},\mathcal{G}_{\mathcal{U}^{{}^{\prime}}}\}$, the generation process is formulated as follows:
+$\text{U}_{t}^{(\cdot)}=f^{\text{gen}}_{\theta}\big{(}p_{\text{gen}},\ \mathcal{U}^{(\cdot)},\ \mathcal{T}_{\mathcal{G}^{(\cdot)}}\big{)},$ (4)
+where $p_{\text{gen}}$ denotes the prompt for generating the response.
+$f^{\text{gen}}_{\theta}$ is the LLM.
+We refer to the generation $\text{U}_{t}$ as TG-DRG-NR, and the generation $\text{U}_{t}^{{}^{\prime}}$ as TG-DRG-R.
+$\mathcal{T}_{\mathcal{G}^{(\cdot)}}$ denotes a textualised graph, formed of a set of text triples.
+Graph Aware Dialogue Response Generation
+Textualised graph-based generation can effectively improve factuality without training, while it is unable to capture the connectivity information within a KG and furthermore affects the factuality of the response.
+To address this limitation, we encode the knowledge graph based on GNNs (Graph Neural Networks) (Kipf 2016).
+To generate responses with a graph, we aim to encode the dialogue sense graph $\mathcal{G}_{\mathcal{U}^{{}^{\prime}}}$ into a graph-level representation.
+We firstly encode its nodes $\mathcal{V}^{{}^{\prime}}$:
+$\mathbf{H}^{0}=\text{GraphEmb}(\mathcal{V}^{\prime})\in\mathbb{R}^{|\mathcal{V}^{\prime}|\times d},$ (5)
+where $d$ is the embedding dimension.
+The matrix $\mathbf{H}^{0}$ consists of node embeddings produced by the Graph Embedding Layer (GraphEmb).
+To get the higher node representation, we utilise a layer based on multi-head attention (Vaswani et al.
+2017), which is formulated as follows:
+$\mathbf{H}^{l+1}=\text{GTC-LN}(\mathbf{H}^{l})\in\mathbb{R}^{|\mathcal{V}^{\prime}|\times d},$ (6)
+where we employ Graph TransformerConv (Shi et al.
+2020) with Layer Norm (GTC-LN) to encode $\mathbf{H}^{0}$ to higher node representations $\mathbf{H}^{l+1}$, which can capture more global structural information and lay the foundation for generating high factually correct responses.
+$\mathbf{h}_{\text{mean}}=\frac{1}{|\mathcal{V}^{\prime}|}\sum_{v\in\mathcal{V}^{\prime}}\mathbf{H}^{l+1}_{v}\in\mathbb{R}^{d},$ (7)
+$\mathbf{h}_{\text{graph}}=\text{GKProj}(\mathbf{h}_{\text{mean}})\in\mathbb{R}^{d}.$ (8)
+Here we mean pooling the node embeddings to get a global-level representation $\mathbf{h}_{\text{mean}}$, and further, a Graph Knowledge Projector (GKProj), implemented by a feed-forward neural network, is used to transfer the $\mathbf{h}_{\text{mean}}$ to a graph-level representation $\mathbf{h}_{\text{graph}}$.
+The projector serves as a bridge between the graph and response generation.
+$\mathbf{z}=\text{TextEmb}(\mathcal{U})\in\mathbb{R}^{d}.$ (9)
+Here we encode the utterances $\mathcal{U}$ to a text embedding $\mathbf{z}$ by the Text Embedding Layer (TextEmb) of an LLM.
+We generate the token $\text{U}^{*}_{ti}$ by sampling the probability of $\text{U}^{*}_{ti}$ conditioned on generated tokens $\text{U}^{*}_{t,<i}$, $\mathbf{h}_{\text{graph}}$, $\mathbf{z}$ and the parameter $\theta$, formulated as follows:
+$\text{U}^{*}_{ti}\sim P(\text{U}^{*}_{ti}\mid\text{U}^{*}_{t,<i},\mathbf{h}_{\text{graph}},\mathbf{z};\theta).$ (10)
+We train the model using the following cross-entropy loss:
+$\mathcal{L}_{\text{CE}}=-\sum_{i=1}^{T}\sum_{c=1}^{C}y_{i}^{(c)}\log p_{i}^{(c)},$ (11)
+where $p_{i}^{(c)}=P(\text{U}^{*}_{ti}=c\mid\text{U}_{t,<i},\mathbf{h}_{\text{graph}},\mathbf{z};\theta)$ is the model-predicted probability over vocabulary class $c$ at step $i$.
+To retain the pretrained knowledge of the LLM, we adopt LoRA combined with cross-entropy to train our graph model.
+We refer the generation $\text{U}^{*}_{t}$ to GA-DRG.
+3.4 Dialogue Fact Score
+We propose the Dialogue Fact Score that incorporates dialogue context and explicitly handles cases where supporting information is unavailable.
+We redesign the atomic fact extraction and verification prompts used in computing FactScore.
+In the redesigned atomic fact extraction prompt, only complete sentences are eligible for splitting, and the model is instructed not to introduce any additional information when generating.
+For fact verification, we revise the prompt to incorporate dialogue history and introduce a new class: “Not Enough Information”.
+If an atomic fact is not a verifiable factual claim or lacks direct support from any known source, the model outputs “Not Enough Information.”
+We leverage the LLaMA 3.3 70B Instruct model to compute the dialogue fact score.
+Since the FactScore does not define the knowledge retrieval process for dialogue, we extract entities from atomic facts and dialogue as Wikipedia titles to retrieve passages as supporting knowledge.
 4 Experiment
 4.1 Dataset
-Two public dialogue datasets, OpendialKG Moon et al.
-(2019) and HybriDialogue Nakamura et al.
-(2022), are used to evaluate our methods.
-The two datasets both provide response references but not factual ones.
-See Appendix B for the details and examples of datasets.
+Two public knowledge-driven and English dialogue datasets, OpendialKG (Moon et al.
+2019) and HybriDialogue (Nakamura et al.
+2022), are used in our work.
+OpenDialKG consists of a recommendation task related to movies and books and a chit-chat task related to sports and music.
+HybriDialogue is an open-domain and information-seeking dialogue dataset.
 4.2 Conventional Evaluation Metrics
 We aim to evaluate factuality and quality for dialogue response generation.
-BLEU Reiter (2018), ROUGE-L Lin (2004), Perplexity Jelinek et al.
-(1977) and F1 score Nan et al.
-(2021) are employed to evaluate the quality and fact consistency of the generated responses.
-For more details, please refer to Appendix E.
+BLEU (Reiter 2018), ROUGE-L (Lin 2004), Perplexity (Jelinek et al.
+1977) and F1 score (Nan et al.
+2021) are employed to evaluate the quality and fact consistency of the generated responses automatically.
+For human evaluation, Coherence checks if the response makes sense and matches the previous context.
+Fluency looks at whether the response is grammatically correct and sounds natural.
+Informativeness measures if the response adds useful or new information to the conversation.
+We randomly extracted 50 samples for each model and recruited two annotators for annotation.
+HybriDialogue achieves Agreement/Kappa scores of 0.83/0.41 for Coherence, 0.90/0.53 for Fluency, and 0.89/0.53 for Informativeness, while OpendialKG scores 0.83/0.47, 0.96/0.27, and 0.80/0.60 on the same metrics, respectively.
+However, Kappa is less effective in reflecting the true agreement under imbalanced conditions.
+In the fluency task of OpendialKG, one class accounts for approximately 97% of the data, making it hard for kappa to accurately measure the agreement.
 4.3 Modified Evaluation Metrics
-We adopt both the adapted fact score and the NEIP (Not Enough Info Proportion) to comprehensively evaluate factuality.
-We also manually assess fact score with humans, as shown in Table 1.
-See the Appendix F for more annotation details.
-Datasets Agreement Cohen’s Kappa
+We adopt both the dialogue fact score and the NEIP (Not Enough Info Proportion) to evaluate factuality comprehensively.
+Two human annotators manually assessed the dialogue fact score with 100 random samples on two datasets, respectively, as shown in Table 1.
+The Cohen’s Kappa scores are both above 0.6, indicating substantial agreement between humans and the dialogue score.
+Datasets Raw Agreement Cohen’s Kappa
 HybriDialogue 0.78 0.65
 OpendialKG 0.78 0.67
-Table 1: The agreement and Cohen’s Kappa between the ground truth and LLM outputs regarding the fact score.
+Table 1: Raw agreement and Cohen’s Kappa between ground truth and LLM dialogue fact scores.
+OpendialKG HybriDialogue
 […]
-ChatGLM-6B 1.42 14.24 12.98 70.91 50.13 15.14
-   +BM25 1.38 14.02 13.15 70.99 49.07 14.92
-   +KAPING 1.31 13.93 13.15 72.37 48.74 15.34
-   +Triton((Ours) 1.37 14.07 13.31 74.64 48.73 15.40
-   +Triton-C(Ours) 1.34 13.47 14.41 76.46 37.93 15.33
-[…]
-   +BM25 2.65 12.90 18.92 69.32 56.09 16.25
-   +KAPING 2.56 12.72 18.18 68.34 55.79 16.03
-   +Triton(Ours) 2.60 12.66 18.89 71.56 54.28 16.09
-   +Triton-C(Ours) 2.19 11.27 20.22 74.25 28.53 15.00
-[…]
-   +BM25 2.68 12.71 23.18 72.06 31.98 13.69
-   +KAPING 2.91 12.82 23.22 73.00 32.07 13.85
-   +Triton(Ours) 3.16 13.06 23.07 76.20 31.42 14.50
-   +Triton-C(Ours) 2.33 10.69 20.03 77.77 20.35 13.78
-[…]
-Triton-X(Ours) 3.99 25.86 14.43 87.73 59.20 13.75
-Table 2: The experimental result with various methods on the OpendialKG dataset.
-The column heads with * indicate our primary metrics for evaluating factual accuracy.
-Ours means generating responses with our proposed methods.
-The bold number is the best result within each model.
-PPL indicates Perplexity and a lower PPL means more fluency.
+ChatGLM-6B 70.91 50.13 15.14 12.98 1.42 14.24 46.58 58.19 39.57 23.66 4.62 24.03
++BM25 70.99 49.07 14.92 13.15 1.38 14.02 53.32 55.18 40.26 23.77 4.68 24.19
++KAPING 72.37 48.74 15.34 13.15 1.31 13.93 55.84 54.51 40.99 23.58 4.56 24.07
++TG-DRG-NR 74.64 48.73 15.40 13.31 1.37 14.07 58.15 53.07 40.96 23.95 4.71 24.43
++TG-DRG-R 76.46 37.93 15.33 14.41 1.34 13.47 59.10 50.99 40.92 26.03 4.53 24.23
+Flan-T5-Large 67.76 54.94 16.35 18.37 2.57 13.00 60.04 50.48 40.04 30.89 9.61 27.12
++BM25 69.32 56.09 16.25 18.92 2.65 12.90 67.42 47.43 38.50 32.79 8.71 25.81
++KAPING 68.34 55.79 16.03 18.18 2.56 12.72 68.92 46.56 40.41 31.96 8.92 26.43
++TG-DRG-NR 71.56 54.28 16.09 18.89 2.60 12.66 70.79 43.52 39.70 30.81 8.90 26.14
++TG-DRG-R 74.25 28.53 15.00 20.22 2.19 11.27 74.41 33.68 40.63 34.01 8.20 24.63
+Flan-T5-XXL 69.29 36.62 14.07 23.37 3.44 13.63 53.06 50.62 44.07 37.47 11.34 29.62
++BM25 72.06 31.98 13.69 23.18 2.68 12.71 60.40 48.79 44.65 38.89 11.08 29.75
++KAPING 73.00 32.07 13.85 23.22 2.91 12.82 63.84 48.94 44.13 37.88 10.95 29.44
++TG-DRG-NR 76.20 31.42 14.50 23.07 3.16 13.06 66.62 46.49 43.91 41.35 10.93 29.30
++TG-DRG-R 77.77 20.35 13.78 20.03 2.33 10.69 69.34 39.23 43.65 32.48 9.89 28.19
+G-Retriever 84.79 62.65 11.70 14.98 3.79 26.00 70.29 55.77 42.05 49.66 20.36 39.53
+GA-DRG 87.73 59.20 13.75 14.43 3.99 25.86 73.00 54.87 43.98 39.12 20.85 40.77
+Table 2: The experimental results on the OpendialKG and HybriDialogue dataset.
+The column heads with * indicate our primary metrics for evaluating factuality.
+Bold models mean generating responses with our proposed methods.
+The bold number is the best result within each category.
 NEIP denotes the “not enough information” proportion, representing the percentage of atomic facts that either lack direct knowledge support or fail to qualify as factual claims.
+PPL indicates Perplexity, and a lower PPL means more fluency.
+Fact, BLEU and ROUGE denote Dialogue Fact Score, BLEU-4 and ROUGE-L, respectively.
 4.4 Baseline Methods
-Several widely recognised LLMs, such as ChatGLM Zeng et al.
-(2022) and Flan-T5 Chung et al.
-(2022), have been selected for this work.
-Furthermore, we compare our proposed Triton, Triton-C and Triton-X with BM25 Robertson et al.
-(2009), KAPING Baek et al.
-(2023), G-Retriever He et al.
-(2024), and a basic dialogue generation model.
-Additional details of LLMs can be found in the Appendix G.
-[…]
-ChatGLM-6B 4.62 24.03 23.66 46.58 58.19 39.57
-   +BM25 4.68 24.19 23.77 53.32 55.18 40.26
-   +KAPING 4.56 24.07 23.58 55.84 54.51 40.99
-   +Triton(Ours) 4.71 24.43 23.95 58.15 53.07 40.96
-   +Triton-C(Ours) 4.53 24.23 26.03 59.10 50.99 40.92
-[…]
-   +BM25 8.71 25.81 32.79 67.42 47.43 38.50
-   +KAPING 8.92 26.43 31.96 68.92 46.56 40.41
-   +Triton(Ours) 8.90 26.14 30.81 70.79 43.52 39.70
-   +Triton-C(Ours) 8.20 24.63 34.01 74.41 33.68 40.63
-[…]
-   +BM25 11.08 29.75 38.89 60.40 48.79 44.65
-   +KAPING 10.95 29.44 37.88 63.84 48.94 44.13
-   +Triton(Ours) 10.93 29.30 41.35 66.62 46.49 43.91
-   +Triton-C(Ours) 9.89 28.19 32.48 69.34 39.23 43.65
-[…]
-Triton-X(Ours) 20.85 40.77 39.12 73.00 54.87 43.98
-Table 3: The experimental result on the HybriDialogue dataset with different baselines. The best results within each model are displayed in bold.
+Several widely recognised LLMs, such as ChatGLM (Zeng et al.
+2022) and Flan-T5 (Chung et al.
+2022), have been selected for this work.
+Furthermore, we compare our proposed TG-DRG-NR, TG-DRG-R and GA-DRG with BM25 (Robertson, Zaragoza et al.
+2009), KAPING (Baek, Aji, and Saffari 2023), G-Retriever (He et al.
+2024).
 4.5 Experimental Setup
-We illustrate the details of the experimental setup in Appendix D.
+We provide detailed information regarding the experimental setup of our framework and the evaluation metrics used.
+Except for the GNN settings where the rank and alpha parameters are set to 8 and 16, respectively, the rank and alpha parameters for LoRA (Hu et al.
+2021) are consistently set to 32.
+As G-retriever is implemented using Llama2 7B (Touvron et al.
+2023) with LoRA, we adopt the same model under identical settings to ensure fairness.
+We utilise the multi-qa-mpnet-base-dot-v1 version for the baseline KAPING.
+All experiments were run a single time.
+We utilised the evaluate package to compute BLEU-4 and ROUGE-L scores. Besides, the PPL is calculated using GPT-2 (Radford et al. 2019).
+We used an A100 GPU for fine-tuning and inference. We utilised two A100 GPUs for the computation of the dialogue fact score. All A100 GPUs we used in this work have 80GB of memory.
 4.6 Experimental Result and Analysis
-Tables 2 and 3 show the experimental results of different methods on two public datasets. We analyze the generated responses from the perspectives of text quality and factual consistency.
+We compare our proposed TG-DRG, which leverages a textualised graph, and GA-DRG, which utilises a GNN-based graph encoding approach, against LLM-based baselines, as well as BM25, KAPING, and G-Retriever.
+Table 2 shows the experimental results of different methods on two public datasets.
+We analyze the generated responses from the perspectives of factuality, text quality and model scale.
+Factuality
+With knowledge augmented methods, the dialogue fact score rises in all LLMs, and our TG-DRG-NR has a higher increase than KAPING and BM25.
+From TG-DRG-R, we can see a continuous improvement in dialogue fact score, which means the dialogue reformulation also contributes a slight improvement to the factuality.
+The NEIP is the proportion of atomic facts that can not be directly found in external knowledge to support or which are not factual claims, and the lower one indicates the responses contain more verifiable facts.
+Compared to baseline LLMs, knowledge-augmented methods generally help reduce this proportion and produce more verifiable facts.
+From the table 2, we observe a consistent decrease in NEIP for the TG-DRG-R compared to KAPING and BM25.
+Furthermore, we noticed a marked improvement in NEIP with TG-DRG-NR.
+This suggests that reformulating the dialogue can substantially enhance factuality.
+Notably, the F1 score is not always consistent with the fact score.
+For instance, the trend of the F1 score shown in Flan-T5-XXL is almost opposite to the dialogue fact score on the OpendialKG dataset.
+The reasons can be attributed to either the reference responses lack explicit answers or the F1 score ignores the semantic aspects, and two sentences with the same entities can express contrary meanings.
+In short, knowledge-augmented methods improve the factuality of LLMs, including the dialogue fact score and NEIP.
+Our proposed frameworks perform better than G-Retriever, KAPING and BM25 in the above metrics.
 Text Quality
 With the KG, the BLEU-4 and ROUGE-L scores fluctuate across two datasets.
-Our proposed Triton, the knowledge retriever, generally perform better than other knowledge-enhanced baselines, while we see a drop with Triton-C, the combination of dialogue rewrite and knowledge retriever.
-Our proposed framework, Triton-X, which consists of three components, achieves the highest BLEU and ROUGE-L scores, surpassing the SOTA G-retriever and demonstrating the superiority of our method.
+Our proposed TG-DRG-NR generally perform better than other knowledge-enhanced baselines, while we see a drop with TG-DRG-R.
+Our proposed GA-DRG achieves the highest BLEU and ROUGE-L scores, surpassing the SOTA G-retriever and demonstrating the superiority of our method.
 There is a slight growth in PPL of smaller language models like ChatGLM and Flan-T5-Large: incorporating the KG decreases text fluency.
-After analysis, we found that the LLMs sometimes can not understand the KG well and thus generate unpolished responses.
-We also utilise Case 1 from Section 4.9 to illustrate the factors contributing to low PPL.
+After analysis, we found that the relatively smaller LLMs sometimes can not understand the KG well and thus generate unpolished responses.
 Overall, knowledge-augmented methods show variable performance across different LLMs in terms of BLEU-4 and ROUGE-L scores, and they tend to increase PPL slightly for smaller language models.
-Our proposed Triton-X gains the best performance in BLEU-4 on both datasets and ROUGE-L on HybriDialogue, indicating a higher quality in generating responses.
-Factuality
-With retrieved methods, the fact score rises in all LLMs, and Triton has a higher increase than KAPING and BM25.
-From Triton-C, we can see a continuous improvement in fact score, which means the rewritten dialogue also contributes a slight improvement to the fact score.
-The NEIP is the proportion of atomic facts that can not be directly found in external knowledge to support or which are not a factual claim, and the lower one indicates the responses contain more verifiable facts.
-Compared to baseline LLMs, knowledge-augmented methods generally help reduce this proportion and produce more verifiable facts.
-From the tables, we observe a consistent decrease in NEIP for the Triton compared to KAPING and BM25.
-Furthermore, we noticed a significant improvement in NEIP with the rewritten dialogue.
-This suggests that rewriting the dialogue can substantially enhance factuality.
-It is worth noting that the F1 score is not always consistent with the fact score.
-For instance, the trend of the F1 score shown in Flan-T5-XXL is almost opposite to the fact score on the OpendialKG dataset.
-One reason is that the reference responses lack explicit answers.
-Another reason is that the F1 score ignores the semantic aspects, and two sentences with the same entities can express contrary meanings.
-In short, knowledge-augmented methods improve the factuality of LLMs, including the fact score and NEIP.
-Our proposed Triton series methods perform better than G-Retriever, KAPING and BM25 in the above metrics.
+Our proposed GA-DRG gains the best performance in BLEU-4 on both datasets and ROUGE-L on HybriDialogue, indicating a higher similarity between responses and references.
 Model Scale
-We conduct experiments with different LLMs. The Flan-T5 series LLMs (Large and XXL) are beneficial for us in analysing the correlation between size and performance.
-As the model scale grows, fact score and NEIP exhibit differing trends across the two datasets—factuality declines on HybriDialogue, whereas the OpendialKG dataset shows the opposite pattern.
-The factuality is not always consistent with the model size, which contradicts our intuition.
-HybriDialogue OpendialKG
-[…]
-w/o Retriever -1.1 -0.11 -1.1 +1.2
-w/o Rewrite -2.3 +0.48 -1.46 +0.44
-w/o Graph -3.41 -0.03 -1.00 -1.52
-Table 4: Ablation study of Triton-X, and w/o means removing corresponding components.
+We conduct experiments with different scales of LLMs. The Flan-T5 series LLMs are beneficial in analysing the correlation between size and performance.
+As model scale grows, dialogue fact score and NEIP show differing trends: factuality declines on HybriDialogue but improves on OpendialKG, which contradicts our intuition that larger models ensure consistent factuality.
+OpendialKG HybriDialogue
+Methods $\downarrow Fact NEIP\downarrow$
+w/o Know. Sel. -1.1 +1.2 -1.1 -0.11
+w/o Dial. Reform. -1.46 +0.44 -2.3 +0.48
+w/o Graph Repr. -1.00 -1.52 -3.41 -0.03
+Table 3: An ablation study on GA-DRG, where “w/o” denotes the removal of components: Knowledge Selection (Know. Sel.), Dialogue Reformulation (Dial. Reform.), and Graph Representation (Graph Repr.).
 4.7 Ablation Study
-We conduct an ablation study with our proposed framework, shown in Table 4.
-We remove our components one by one from Triton-X to observe the performance.
-We only report factually related metrics as they are our aim.
-The decline in performance after removing either the rewrite or the retriever component highlights their importance.
+We conduct an ablation study with GA-DRG, shown in Table 3.
+We remove our components one by one from GA-DRG to observe the performance.
+We only report our aim—factually related metrics.
+The results show a noticeable decline in performance after removing each component, underscoring their critical roles in maintaining the factual accuracy of GA-DRG.
 4.8 Human Evaluation Results
-HybriDialogue OpendialKG
+OpendialKG HybriDialogue
 Methods Cohe. Flu. Info. Cohe. Flu. Info.
-ChatGLM-6B 1.90 1.99 1.71 1.99 1.91 1.11
-+bm25 1.85 1.97 1.75 1.99 1.89 1.09
-+SenEmb 1.84 1.98 1.71 1.98 1.95 1.18
-+Triton 1.93 1.98 1.76 2.00 1.94 1.11
-+Triton-C 1.94 1.96 1.85 1.98 1.97 1.28
-Flan-T5-XXL 1.62 1.74 1.74 1.56 2.00 1.36
-+bm25 1.67 1.72 1.80 1.38 1.94 1.49
-+SenEmb 1.59 1.80 1.77 1.46 1.99 1.53
-+Triton 1.59 1.81 1.77 1.52 1.98 1.39
-+Triton-C 1.51 1.83 1.76 1.47 1.97 1.65
-G-retriever 1.92 1.82 1.84 1.96 1.96 0.71
-TritonX 1.96 1.81 1.82 2.00 2.00 0.60
+ChatGLM-6B 1.99 1.91 1.11 1.90 1.99 1.71
++BM25 1.99 1.89 1.09 1.85 1.97 1.75
++KAPING 1.98 1.95 1.18 1.84 1.98 1.71
++TG-DRG-NR 2.00 1.94 1.11 1.93 1.98 1.76
++TG-DRG-R 1.98 1.97 1.28 1.94 1.96 1.85
+Flan-T5-XXL 1.56 2.00 1.36 1.66 1.69 1.74
++BM25 1.38 1.94 1.49 1.68 1.71 1.80
++KAPING 1.46 1.99 1.53 1.58 1.78 1.77
++TG-DRG-NR 1.52 1.98 1.39 1.62 1.76 1.77
++TG-DRG-R 1.47 1.97 1.65 1.55 1.85 1.76
+G-retriever 1.96 1.96 0.71 1.92 1.82 1.84
+GA-DRG 2.00 2.00 0.60 1.96 1.81 1.82
 Table 4: Human evaluation results for Coherence (Cohe.), Fluency (Flu.), and Informativeness (Info.). The best results are marked within each category.
-We report the human evaluation with the perspective: coherence, fluency and informativeness in Table 5.
-Each aspect is rated as one of three levels: positive, neutral, or negative (e.g., coherent, neutral, incoherent).
-Our proposed Triton series methods demonstrate notable improvements in coherence across various settings, with the exception of the Flan-T5-XXL model.
+We present the human evaluation results from three perspectives: coherence, fluency, and informativeness, as shown in Table 4.
+Each aspect is assessed on a three-level scale: positive (2 points), neutral (1 point), or negative (0 points) (e.g., coherent, neutral, incoherent).
+The average scores for each aspect are then calculated based on these ratings.
+Our proposed frameworks TG-DRG and GA-DRG demonstrate marked improvements in coherence across various settings, with the exception of the Flan-T5-XXL model.
 In terms of fluency, our methods achieve the highest scores for Flan-T5-XXL on HybriDialogue, as well as for ChatGLM and graph-based methods on OpendialKG.
 Overall, the proposed methods maintain strong performance in fluency and coherence.
-Additionally, the Triton series generally enhances informativeness.
+Additionally, the TG-DRG generally enhance informativeness.
+Case 1
+Dialogue {Original / Coreference Resolved}:
+Speaker A: What do you think about Wladimir Klitschko?
+Speaker B: Not a whole lot but I do know {he / Wladimir Klitschko} is married to Hayden Panettiere from the show Nashville.
+Speaker A: I have watched {that / Nashville} before, is {she / Hayden Panettiere} in any other show?
+Speaker B: {She / Hayden Panettiere} played a voice on the cartoon A Bug’s Life. Other than that I can’t think of where I’ve seen {her / Hayden Panettiere}.
+Speaker A: Ok, I didn’t know {she / Hayden Panettiere} voiced a character in {that / A Bug’s Life}.
+Reference: I was surprised to hear that Kevin Spacey also voiced a character in that movie.
+Flan-T5-XXL: She voiced a character in the cartoon A Bug’s Life.
+BLEU-4: 13.04; ROUGE-L: 30.8
+Key Knowledge: (Dot, voice actor, Hayden Panettiere)
+TG-DRG-R: Hayden Panettiere voiced the character Dot in A Bug’s Life.
+BLEU-4: 0; ROUGE-L: 23.1
+Table 5: Dialogue and LLMs’ responses related to Hayden Panettiere, before and after coreference resolution.
 4.9 Case Study
-We use two cases that are beneficial for us in understanding our methods’ limitations. See more descriptions in Appendix H.
+We use a case about Hayden Panettiere shown in Table 5 to illustrate our work.
+From the case, the dialogue contains pronouns, e.g.
+“she”, challenging the generation of factually correct responses.
+We can see the response from Flan-T5-XXL is repeating the previous utterance “She played a voice on the cartoon A Bug’s Life…”.
+After dialogue reformulation and dialogue sense knowledge selection, our TG-DRG-R is able to generate a more factual and detailed response.
+However, from Table 2, we observe that TG-DRG-R’s response achieves lower BLEU-4 and ROUGE-L scores compared to those of Flan-T5-XXL.
+This suggests that BLEU-4 and ROUGE-L may not effectively capture the quality of responses in open-domain dialogue tasks.
 5 Conclusion
-Given the limitations of previous methods in accurately retrieving knowledge triples and their challenges in handling dialogue tasks due to coreference issues, we propose a novel framework.
-This framework features a carefully designed knowledge retriever to identify valuable knowledge triples, an advanced dialogue rewriter to resolve coreference effectively, and a knowledge-enhanced dialogue generation component to improve factuality.
-Building on this, we also adapt the fact score to make it suitable for evaluating the factuality of the dialogue response and verifying its validity in the dialogue task.
-The results demonstrate that our method significantly enhances the factual accuracy of dialogue responses compared to other baselines, including the SOTA G-retriever.
-We analyse some cases and see some limitations in our current work. We plan to explore the mechanism of LLM reactions to triples in the future.
-Limitations
-Although we propose a well-designed framework for dialogue-related tasks, some limitations remain.
-Our current knowledge retrieval method relies on the exact entity overlap between the query and Wikidata.
-This approach lacks semantic understanding, making it inefficient in retrieving semantically related knowledge triples.
-Constructing a graph database from Wikidata could be an effective way to address this problem, but building such a large-scale database presents significant challenges.
-Additionally, our approach relies on knowledge triples.
-However, the validity of certain triples, such as (“Montevideo”, “population”, “309331”) shown in the key knowledge of Case 1 in Section 4.9, will change over time.
-Retrieving such triples for LLMs could lead to outdated responses.
-Ethical Statement
-In this work, we follow the ethical standards in data use and research.
-The datasets used in this study, OpendialKG and HybriDialogue, are publicly available and have been widely used in research.
-We have not added any extra personal information to these publicly available datasets.
-The data from human annotators was limited to some specific labels, and the results reported in this paper are based on statistical analysis.
-No personal or sensitive information was accessed during the research process.
+Our work addresses the limitations of previous methods in handling factual dialogue response generation, particularly issues arising from complex coreference in the dialogue.
+To this end, we propose two novel frameworks, TG-DRG and GA-DRG, which feature reasoning-guided dialogue reformulation and dialogue sense knowledge selection, leveraging graph-based generation techniques.
+The results demonstrate that these frameworks noticeably enhance the factual accuracy of dialogue responses compared to baseline methods, including the SOTA G-retriever.
+Additionally, we propose the dialogue fact score as a novel metric to evaluate the factuality of dialogue responses.
+This metric shows high consistency with human evaluations and proves to be more reliable than traditional metrics such as BLEU-4, ROUGE-L, and F1, which often fail to capture factual consistency in open-domain dialogue tasks.
+While our approach shows noticeable improvements, certain limitations remain.
+For instance, the validity of triples from external knowledge graphs may change over time, potentially impacting factuality.
+Addressing this challenge will be a focus of our future work.
 Appendix A Prompts
-The prompts of the knowledge-augmented approach are shown in Table 6. The factuality-related prompts are described in Table 8.
-Dialogue Rewrite Prompt
+In this section, we show the prompts used in our work.
+The prompt for reasoning-guided dialogue reformulation is described in Table 6.
+The prompts for the knowledge-augmented dialogue response generation, used in TG-DRG, and atomic fact splitting are shown in Table 7.
+The prompt for fact verification of dialogue fact score prompts is described in Table 8.
+Dialogue Reformulation Prompt
 You are tasked with resolving all pronouns and references in the given dialogue to their explicit entities.
@@
 Output Format: **Chain of Thought**: [Your reasoning process for resolving the references] **Resolved Dialogue**: [The dialogue with all pronouns and references resolved] Dialogue: {Dialogue}
-Table 6: Prompt for rewriting the dialogue and generating a response with dialogue and knowledge triples.
+Table 6: The prompt for reasoning-guided dialogue reformulation.
 Dialogue Response Generation Prompt Atomic Fact Splitting Prompt
 Knowledge: {Selected Triples} Dialogue: {Dialogue Context} Given the above knowledge and dialogue, please respond to the input below and ensure the response is fluent and fact-consistent in English.
-Input: {Query} Response: … {Examples..} If the following input is an incomplete sentence or a phrase, please output it exactly as it is.
+Input: {Last Utterance} Response: … {Examples..} If the following input is an incomplete sentence or a phrase, please output it exactly as it is.
 Otherwise, if it is a complete sentence, split it into atomic sentences based only on the given information, without adding any additional information or making inferences: Input: {Response} Output …:
-Table 7: The prompts for getting atomic facts and fact score.
-Fact Score Prompt
+Table 7: The prompts for dialogue response generation (left) and atomic fact splitting (right).
+Dialogue Fact Score Prompt
 Instruction: The statement is part of a response in a dialogue.
 Evaluate the statement strictly based on the provided knowledge source and dialogue history only.
-If the statement is not a factual claim (e.g., opinion, question, or unclear assertion), output: "no enough information." If it is a factual claim: Output true if the statement is directly supported by evidence in the knowledge source or dialogue history.
+If the statement is not a factual claim (e.g., opinion, question, or unclear assertion), output: ”no enough information.” If it is a factual claim: Output true if the statement is directly supported by evidence in the knowledge source or dialogue history.
 Output false if the statement is directly contradicted by the knowledge source or dialogue history.
@@
 Evidence: {Wikipedia Passages} Dialogue history: {Dialogue} Speaker A: {Speaker} Statement: {Atomic Fact}
-Table 8: The prompts for getting atomic facts and fact score.
+Table 8: The prompt for fact verification of dialogue fact score.
 Appendix B Datasets Details
@@
 OpenDialKG
-It is an English knowledge-driven dialogue dataset that consists of a recommendation task related to movies and books and a chit-chat task related to sports and music.
+It is a knowledge-driven dialogue dataset.
 The dataset contains 13,802 samples.
@@
 Finally, we randomly selected 15% (1,962 samples) for the validation set, another 15% (1,973 samples) for the test set, and the remaining 70% (9,120 samples) for the training set.
-The dataset provides a number of triples extracted from Freebase Bast et al.
-(2014).
+The dataset provides a number of triples extracted from Freebase (Bast et al.
+2014).
 These triples were collected several years ago.
 To address any potential issues with outdated information, we re-extract the triples from Wikidata.
-OpendialKG is released under CC-BY-NC-4.0 license [^2] https://github.com/facebookresearch/opendialkg, which permits non-commercial research use.
+OpendialKG is released under CC-BY-NC-4.0 license, which permits non-commercial research use.
 HybriDialogue
-It is an open-domain and information-seeking dialogue dataset in English.
+It is an open-domain dialogue dataset.
 It is constructed by splitting complex questions into multi-turn dialogue.
@@
 The training set contains 4,359 samples, while the validation set includes 242 samples, and the test set consists of 243 samples.
-HybriDial follows the MIT license Nakamura et al.
-(2022), which allows both commercial and non-commercial use.
+HybriDial follows the MIT license (Nakamura et al.
+2022), which allows both commercial and non-commercial use.
 OpendialKG Example
-Query: Could you recommend some movies staring Chiaki Kuriyama?
-Retrieved triples from BM25: (Chiaki Kuriyama, given name, Chiaki) (Chiaki Kuriyama, family name, Kuriyama) (Chiaki Kuriyama, place of birth, Tsuchiura) (Chiaki Kuriyama, occupation, film actor) (Chiaki Kuriyama, instrument, voice) …
-Retrieved triples from KAPING: (Chiaki Kuriyama, given name, Chiaki) (Chiaki Kuriyama, occupation, model) (Chiaki Kuriyama, occupation, singer) (Chiaki Kuriyama, occupation, actor) (Chiaki Kuriyama, occupation, fashion model) …
-Retrieved triples from Triton: (Kill Bill Volume 1, cast member, Chiaki Kuriyama), (Into the Sun, cast member, Chiaki Kuriyama), (Kagen no Tsuki, cast member, Chiaki Kuriyama), (Kamogawa Horumo, cast member, Chiaki Kuriyama), (Gonin, cast member, Chiaki Kuriyama) …
+Last Utterance: Could you recommend some movies staring Chiaki Kuriyama?
+Selected triples from BM25: (Chiaki Kuriyama, given name, Chiaki) (Chiaki Kuriyama, family name, Kuriyama) (Chiaki Kuriyama, place of birth, Tsuchiura) (Chiaki Kuriyama, occupation, film actor) (Chiaki Kuriyama, instrument, voice) …
+Selected triples from KAPING: (Chiaki Kuriyama, given name, Chiaki) (Chiaki Kuriyama, occupation, model) (Chiaki Kuriyama, occupation, singer) (Chiaki Kuriyama, occupation, actor) (Chiaki Kuriyama, occupation, fashion model) …
+Selected triples from our method: (Kill Bill Volume 1, cast member, Chiaki Kuriyama), (Into the Sun, cast member, Chiaki Kuriyama), (Kagen no Tsuki, cast member, Chiaki Kuriyama), (Kamogawa Horumo, cast member, Chiaki Kuriyama), (Gonin, cast member, Chiaki Kuriyama) …
 HybriDialogue Example
-Query: Hi. Can you tell me who Judd Trump is?
-Retrieved triples from BM25: (Judd Trump, given name, Judd)(Judd Trump, family name, Trump)(2021 Champion of Champions, winner, Judd Trump)(Judd Trump, victory, 2016 China Open)(Judd Trump, victory, 2011 China Open) …
-Retrieved triples from KAPING: (Judd Trump, occupation, snooker player) (Judd Trump, nickname, The Ace) (Judd Trump, given name, Judd) (Judd Trump, family name, Trump) (Judd Trump, place of birth, Whitchurch) …
-Retrieved triples from Triton: (2019 World Snooker Championship, winner, Judd Trump), (Judd Trump, award received, Snooker Hall of Fame), (Judd Trump, country of citizenship, United Kingdom), (Judd Trump, sport, snooker), (Judd Trump, award received, player of the year award) …
+Last Utterance: Hi. Can you tell me who Judd Trump is?
+Selected triples from BM25: (Judd Trump, given name, Judd)(Judd Trump, family name, Trump)(2021 Champion of Champions, winner, Judd Trump)(Judd Trump, victory, 2016 China Open)(Judd Trump, victory, 2011 China Open) …
+Selected triples from KAPING: (Judd Trump, occupation, snooker player) (Judd Trump, nickname, The Ace) (Judd Trump, given name, Judd) (Judd Trump, family name, Trump) (Judd Trump, place of birth, Whitchurch) …
+Selected triples from our method: (2019 World Snooker Championship, winner, Judd Trump), (Judd Trump, award received, Snooker Hall of Fame), (Judd Trump, country of citizenship, United Kingdom), (Judd Trump, sport, snooker), (Judd Trump, award received, player of the year award) …
 Table 9: The examples extracted from OpendialKG and HybriDialogue dataset. The triples are retrieved from Wikidata and selected using different methods.
 Table 9 presents the two examples of OpendialKG and HybridDial datasets with selected triples from different methods.
-Appendix C Examples from GPT-4o
-We employ GPT-4o to annotate query-triple pairs and resolve dialogue coreference from the training dataset.
-In the end, we collected 29K samples of query-triple pairs and 35.4K samples of coreference resolution with reasoning steps.
+Appendix C Examples generated from GPT-4o
+We employ GPT-4o to generate data for dialogue sense knowledge selection and dialogue reformulation from the training dataset, where we collected 29K samples and 35.4K for them.
 The examples for fine-tuning are listed in Tables 10 and  11.
-Query Triple Relevance
+Last Utterance Triple Relevance
 Which teams has the Colorado Avalanche played against in the playoffs? (Colorado Avalanche, sport, ice hockey) Irrelevant
@@
 Hi, do you by any chance know anything about Vladimir Horowitz? (Stefanos Korkolis, student of, Vladimir Horowitz) Relevant
-Table 10: The query and triple are sourced from the training dataset and Wikidata. The relevance is annotated by GPT-4o.
+Table 10: The dialogue and corresponding triple are taken from the training dataset and Wikidata, respectively, with only the last utterance of the dialogue used here..
+The relevance is generated by GPT-4o.
+These examples are adopted for fine-tuning our dialogue sense knowledge selection.
 Section Content
@@
 Speaker A: Oh, I think I’ve read The Magicians it a High Fantasy book.
-CoT Speaker A mentions "Lev Grossman" as the author they like.
-This establishes that "his" in their request for book recommendations refers to Lev Grossman.
-Speaker B mentions "He wrote The Magician’s Land and The Magicians." The pronoun "He" refers to Lev Grossman, as he is the author being discussed.
-Speaker A uses "it" in "I think I’ve read The Magicians it a High Fantasy book." The pronoun "it" refers to "The Magicians," as it is the last explicitly mentioned book in the conversation.
-Resolved Dialogue Speaker A: I like Lev Grossman books can you recommend some of Lev Grossman’s books Speaker B: Sure!
+CoT Speaker A mentions ”Lev Grossman” as the author they like.
+This establishes that ”his” in their request for book recommendations refers to Lev Grossman.
+Speaker B mentions ”He wrote The Magician’s Land and The Magicians.” The pronoun ”He” refers to Lev Grossman, as he is the author being discussed.
+Speaker A uses ”it” in ”I think I’ve read The Magicians it a High Fantasy book.” The pronoun ”it” refers to ”The Magicians,” as it is the last explicitly mentioned book in the conversation.
+Reformulated Dialogue Speaker A: I like Lev Grossman books can you recommend some of Lev Grossman’s books Speaker B: Sure!
 Lev Grossman wrote The Magician’s Land and The Magicians.
@@
 The Magicians is a High Fantasy book.
-Table 11: An example generated by GPT-4o illustrates how CoT reasoning works in coreference resolution.
-Appendix D Experimental Setup
-We provide detailed information regarding the experimental setup of our framework and the evaluation metrics used.
-For fine-tuning the Triton, we only collect the query-triple pairs from the HybriDial dataset.
-We fine-tuned 3 epochs for either the LoRA method or the full-tuning method.
-The rank and alpha parameters for LoRA are both set to 8.
-We utilise the multi-qa-mpnet-base-dot-v1 version for the baseline KAPING.
-As G-retriever is implemented using Llama2 7B Touvron et al.
-(2023), we adopt the same model in our Triton-X under identical settings to ensure fairness.
-For knowledge-augmented methods, the maximum number of selected triples is set to 5.
-Additionally, all experiments were run a single time.
-Regarding the evaluation metrics, we utilised the evaluate package  [^3] https://huggingface.co/docs/evaluate to compute BLEU-4 and ROUGE-L scores.
-Besides, the PPL is calculated using GPT-2 Radford et al.
-(2019).
-We fine-tune our relevance classifier using a single A100 GPU and our dialogue generation model with three A100 GPUs.
-We utilized two A100 GPUs for inference tasks.
-All A100 GPUs we used in this work have 80GB of memory.
-In total, our GPU runtime amounted to approximately 575 hours.
-Appendix E Conventional Evaluation Metrics in Detail
+Table 11: An example generated by GPT-4o illustrates how CoT reasoning works in dialogue reformulation.
+Example 1
+Dialogue:
+Speaker A: I enjoy Charlaine Harris’s work, can you name something similar?
+Speaker B: The Southern Vampire Mysteries I would recommend for you.
+Speaker A: When was it released?
+Speaker B: 2004.
+Speaker A / Reformulated
+How many books are in the series? / How many books are in The Southern Vampire Mysteries series?
+Label: Good
+Reason: The reference “The series” in the last utterances has been resolved as “The Southern Vampire Mysteries” correctly.
+Example 2
+Speaker A: Could you recommend a book written by author Fannie Flagg?
+Speaker B: Welcome to the World, Baby Girl and I Still Dream About You are great books that I would suggest. You read them?
+Speaker A: No, I have read neither of those. What are their genres?
+Speaker B: It is adventure and is also a movie released in 2010, I would suggest reading the book then watching the movie.
+Speaker A / Reformulated:
+Yes, I will definitely read the book first. Can you add it to my list? / Yes, I will definitely read “Welcome to the World, Baby Girl” first. Can you add “Welcome to the World, Baby Girl” to my list?
+Label: Neutral
+Reason: The reference “the book” in the dialogue is ambiguous. So the reformulated utterance may not be correct.
+Example 3
+Dialogue:
+Speaker A: Cool! do you know the occasion name of the Video event?
+Speaker B: Yes, it is NeXT’s internal marketing strategy video.
+Speaker A: Great! do you know the city where that video took place?
+Speaker B: Yes, it was Redwood City.
+Speaker A / Reformulated:
+Good! do you know how many miles that city far from south of San Francisco? / Good! do you know how many miles that city far from south of San Francisco?
+Label: Bad
+Reason: The reference “that city” should be resolved, but the model fails to resolve it.
+Table 12: The manual annotated examples for dialogue reformulation. In human evaluation, we only evaluate the last utterance.
+Appendix D Conventional Evaluation Metrics in Detail
 We list all the details of conventional evaluation metrics as follows:
-  • BLEU Reiter (2018), used in the machine translation initially, calculates the n-gram overlap between generated text and references and reflects the correlation between generated text and human writing.
-  • ROUGE-L Lin (2004) is calculated based on the length of the longest common subsequence, which captures the word order and sentence-level structure from the ground truth.
-  • PPL Jelinek et al. (1977) is widely used in measuring text fluency of generated output, in which calculation is based on the language model.
-  • F1 Score Nan et al. (2021) is based on the precision and recall of the extracted entities matching between the generated response and ground truth.
-Case 1 Case 2
-Context: A: What is the capital and largest city in Uruguay?
-B: Montevideo is the capital and largest city in Uruguay.
-Query: What was the population of Montevideo, Uruguay, in 2011?
-Response: The population of Montevideo, Uruguay, in 2011 was 59,027.
-Key Knowledge: (Montevideo, population, 309331) New Response: Montevideo, population, 309331 Dialogue: A: What kind of game was “Team Fortress 2”?
-B: It’s a multiplayer first-person shooter.
-A: And what was it’s release date?
-Response: It was released on October 26, 2007.
-Rewritten Dialogue: B: What kind of game was “Team Fortress 2”?
-A: “Team Fortress 2” is a multiplayer first-person shooter.
-A: And what was “Team Fortress 2”’s release date?
-New Response: “Team Fortress 2” was released on November 20, 2007.
-Table 12: The responses of cases were both generated by Flan-T5-Large. The new response was created using key knowledge or a rewritten dialogue. Triton selected the key knowledge.
-Appendix F Details of Manual Annotation
-We invited some students from Asian countries to annotate samples.
+  • BLEU (Reiter 2018), used in the machine translation initially, calculates the n-gram overlap between generated text and references and reflects the correlation between generated text and human writing.
+  • ROUGE-L (Lin 2004) is calculated based on the length of the longest common subsequence, which captures the word order and sentence-level structure from the ground truth.
+  • PPL (Jelinek et al. 1977) is widely used in measuring text fluency of generated output, in which calculation is based on the language model.
+  • F1 Score (Nan et al. 2021) is based on the precision and recall of the extracted entities matching between the generated response and ground truth.
+Case 2
+Original Dialogue:
+Speaker A: Hello! do you know fictional marsupials?
+Speaker B: Yes, it is subsidiary to the list of fictional animals and is a collection of various notable marsupial characters that appear in various works of fiction.
+Speaker A: Great! can you list some literature names of fictional marsupials?
+Speaker B: Yes, there are many books available including Ginger Kangaroo, Kidding Kangaroo, etc.
+Speaker A: Good! do you know the species name of the literature Blinky Bill?
+Speaker B: Yes, it is Koala.
+Speaker A: Great! do you know the author of that book?
+Flan-T5-XXL: Yes, it is written by E. Nesbit.
+Fact Verification: False
+Reformulated Dialogue:
+Speaker A: Hello! do you know fictional marsupials?
+Speaker B: Yes, fictional marsupials are subsidiary to the list of fictional animals and is a collection of various notable marsupial characters that appear in various works of fiction.
+Speaker A: Great! can you list some literature names of fictional marsupials?
+Speaker B: Yes, there are many books available including Ginger Kangaroo, Kidding Kangaroo, etc.
+Speaker A: Good! do you know the species name of the literature Blinky Bill?
+Speaker B: Yes, the species name of Blinky Bill is Koala.
+Speaker A: Great! do you know the author of Blinky Bill?
+Key Knowledge: (Blinky Bill, illustrator, Dorothy Wall), (Blinky Bill, author, Dorothy Wall) …
+TG-DRG-R: Yes, the author of Blinky Bill is Dorothy Wall.
+Fact Verification: True
+Case 3
+Context:
+Speaker A: What is the capital and largest city in Uruguay?
+Speaker B: Montevideo is the capital and largest city in Uruguay.
+Speaker A: What was the population of Montevideo, Uruguay, in 2011?
+Flan-T5-Large: The population of Montevideo, Uruguay, in 2011 was 59,027.
+Key Knowledge: (Montevideo, population, 309331)
+TG-DRG-NR: Montevideo, population, 309331
+Table 13: The responses of cases were generated by Flan-T5 series models. In case 2, the resolved coreference is bold in the reformulated dialogue.
+Appendix E Details of Manual Annotation
+We invited several annotators to annotate the samples.
 All of them are well-educated and possess good English proficiency.
 The annotators were told that the annotated data would be used for research, and we got consent for data use.
-To assess the accuracy of query-triple pair annotated from GPT-4o, we invite tree annotators to this task.
-We randomly selected 100 samples and instructed participants to determine whether the query and triple were relevant.
+To assess the accuracy of the samples used in fine-tuning dialogue sense knowledge selection generated from GPT-4o, we invite three annotators to this task.
+We randomly selected 100 samples and instructed participants to determine whether the last utterance of the dialogue is relevant to the triple.
 The output is either relevant or irrelevant.
@@
 These results indicate a high level of accuracy.
-Three annotators were involved in the assessment of the fact score task.
-We used the fact score prompt, described in Table 8, as an instruction.
-Two annotators independently assessed the fact score for 100 samples from the HybriDial and OpendialKG datasets respectively.
+Three annotators were involved in the assessment of the dialogue fact score task.
+We used the dialogue fact score prompt, described in Table 8, as an instruction.
+Two annotators independently assessed the dialogue fact score for 100 samples from the HybriDial and OpendialKG datasets, respectively.
 Initially, the raw agreement scores were 0.76 for HybriDial and 0.81 for OpenDialKG.
@@
 Finally, we report the final agreement and Cohen’s Kappa score with the evaluation model, as shown in Table 1.
-Two annotators were involved in the dialogue human evaluation.
-The accuracy of human annotations is consistently high across all aspects, with HybriDial achieving 0.81 for coherence, 0.88 for fluency, and 0.89 for informativeness, while OpendialKG reaches 0.83, 0.96, and 0.80 respectively, indicating the overall reliability of manual evaluations.
-We finally report the average score in our dialogue human evaluation.
-Appendix G Baseline Descriptions
+We randomly selected 100 samples to evaluate the quality of dialogue reformulation, with two annotators involved in the process.
+The evaluation focuses solely on the last utterance, as it is the most crucial part of the dialogue.
+Each sample is manually rated as good, neutral, or bad.
+On the HybriDialogue dataset, the annotators achieved a raw agreement of 0.91 and a Cohen’s kappa of 0.580.
+On the OpenDialKG dataset, they reached a raw agreement of 0.94 and a Cohen’s kappa of 0.548, indicating moderate agreement.
+The annotation examples of dialogue reformulation are shown in Table 12.
+Appendix F Baseline Descriptions
 LLMs used in this work are presented as follows:
-  • ChatGLM Zeng et al.
-(2022) is an open-source and bilingual language model, based on GLM Du et al.
-(2021) architecture.
+  • ChatGLM (Zeng et al.
+2022) is an open-source and bilingual language model, based on GLM (Du et al.
+2021) architecture.
 The technique of ChatGLM-6B is similar to ChatGPT, optimized in dialogue tasks.
-  • Flan-T5 Chung et al.
-(2022) proposed several instruction-based LLMs, scaling the number of tasks and model size and fine-tuning in the chain-of-thought data.
+  • Flan-T5 (Chung et al.
+2022) proposed several instruction-based LLMs, scaling the number of tasks and model size and fine-tuning in the chain-of-thought data.
 We adopt Flan-T5-XXl (11B), Flan-T5-Large (780M) and Flan-T5-Small (80M) as baselines in our work.
@@
   • No Knowledge: We feed the prompt into LLMs to generate dialogue responses without external knowledge.
-  • BM25 Robertson et al.
-(2009): It is the ranking function between a query and several documents widely used in information retrieval.
+  • BM25 (Robertson, Zaragoza et al.
+2009): It is the ranking function between a query and several documents widely used in information retrieval.
 In our experiments, we also apply RefinED to retrieve triples for BM25 and replace documents with triples for retrieval.
-  • KAPING Baek et al.
-(2023): It is a popular knowledge-augmented method for question answering.
+  • KAPING (Baek, Aji, and Saffari 2023): It is a popular knowledge-augmented method for question answering.
 Since the authors have not released the official code, we follow the experimental setup of the KAPING method: we first extract entities from the query using REfinED, rank the query with triples using MPNet, select the top-N most relevant triples, and supplement them as prompts to generate dialogue responses.
-  • G-Retriever He et al.
-(2024): It is a retrieval-augmented generation method for textual graphs that selects informative subgraphs via a Prize-Collecting Steiner Tree, encodes them with GNNs, and generates responses with high accuracy, outperforming strong baselines and reducing hallucinations.
-  • Triton(Ours): We use GPT-4o to generate high-quality query-triple pairs based on the training dataset and fine-tune Llama3 8B base model with collected samples.
-The model acts as a triple retriever to help select valuable triples for generating dialogue responses.
-  • Triton-C(Ours): We fine-tune a coreference resolution model based on Llama 8B base model using reasoning distillation.
-Our approach integrates dialogue rewriting with entity linking for enhanced performance.
-  • Triton-X(Ours): It is the framework that consists of our well-designed knowledge retriever, dialogue rewrite and graph-based module.
-Appendix H Details of the Case Study
-The examples of case studies are shown in Table 12.
-The topic in Case 1 is about the population of Montevideo, and our methods do help select the relevant triples.
-However, the LLM then generates one of triples directly, which is not fluent and has a higher PPL i.e.
+  • G-Retriever (He et al.
+2024): It is a retrieval-augmented generation method for textual graphs that selects informative subgraphs via a Prize-Collecting Steiner Tree, encodes them with graph neural networks, and generates responses with high accuracy, outperforming strong baselines and reducing hallucinations.
+Appendix G Dialogue Reformulation Evaluation
+We evaluate the dialogue reformulation performance through human judgment, with annotation details provided in  E.
+Each instance is rated as good (1), neutral (0.5), or bad (0), and the final score is computed as the average.
+Our method achieves a score of 0.9325 on the HybriDialogue dataset and 0.9475 on the OpenDialKG dataset, demonstrating the strong effectiveness of our proposed reasoning-guided dialogue reformulation approach.
+Appendix H Addtional Case Studies
+The additional case studies are shown in Table 13.
+Case 2 presents a dialogue about the book Blinky Bill.
+The original dialogue contains coreference issues, making the dialogue difficult to follow and resulting in a factually incorrect response.
+By applying our TGDRG-R approach, which combines step-by-step reference resolution in reasoning-guided dialogue reformulation with dialogue sense knowledge selection, we achieve the generation of correct and coherent responses.
+The topic in Case 3 is about the population of Montevideo, and our methods do help select the relevant triples.
+However, the TG-DRG-NR then generates one of the triples directly, which is not fluent and has a higher PPL i.e.
 lower quality.
-Case 2 is a dialogue that asks about the release date of “Team Fortress 2”, and the original response responds to the dialogue and is close to the writing of reference.
-After rewriting the dialogue, the LLM generates a response that differs from the reference, causing a lower BLEU-4 score.